### v1

# UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers

Abstract
Neural NLP models are often miscalibrated, assigning high confidence to incorrect predictions, which undermines selective prediction and high-stakes deployment. Post-hoc calibration methods adjust output probabilities but leave internal computation unchanged, while ensemble and Bayesian approaches improve uncertainty at substantial training or storage cost. We propose UAT-Lite, an inference-time framework that makes self-attention uncertainty-aware using approximate Bayesian inference via Monte Carlo dropout in pretrained transformer classifiers. Token-level epistemic uncertainty is estimated from stochastic forward passes and used to modulate self-attention during contextualization, without modifying pretrained weights or training objectives. We additionally introduce a layer-wise variance decomposition to diagnose how predictive uncertainty accumulates across transformer depth. Across the SQuAD 2.0 answerability, MNLI, and SST-2, UAT-Lite reduces Expected Calibration Error by approximately 20% on average relative to a fine-tuned BERT-base baseline while preserving task accuracy, and improves selective prediction and robustness under distribution shift. 111github.com/eliashossain001/uq_decomposition
UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers
Elias Hossain1††thanks: Corresponding author: mdelias.hossain@ucf.edu, Shubhashis Roy Dipta2, Subash Neupane3, Rajib Rana4, Ravid Shwartz-Ziv5, Ivan Garibay1, Niloofar Yousefi1 1University of Central Florida 2University of Maryland, Baltimore County 3Meharry Medical College 4University of Southern Queensland 5New York University mdelias.hossain@ucf.edu
1 Introduction
Pretrained transformer-based language models have achieved state-of-the-art performance across a wide range of NLP tasks, from question answering to clinical decision support Devlin et al. (2019); Brown et al. (2020). Despite these advances, modern transformers exhibit systematic miscalibration, frequently assigning high confidence to incorrect predictions Desai and Durrett (2020). This misalignment between confidence and correctness is particularly problematic for selective prediction and related high-stakes applications.
Fig.˜1 summarizes existing approaches to uncertainty estimation. Standard transformer inference is deterministic and does not represent uncertainty during internal computation (Fig.˜1A). Post-hoc calibration methods, such as temperature scaling, adjust confidence only after representations are formed, leaving internal reasoning unchanged (Fig.˜1B; Guo et al., 2017). Stochastic approaches, including Monte Carlo (MC) dropout and deep ensembles (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017), improve uncertainty estimates but largely treat uncertainty as an output-level signal, without influencing attention.
This limitation reflects a broader gap: transformers are trained to optimize predictive accuracy rather than confidence reliability Guo et al. (2017). While post-hoc methods are lightweight and effective in-domain, they do not alter token interactions. Bayesian approaches integrate uncertainty more deeply but typically require architectural modification or specialized training, limiting compatibility with pretrained models. This raises the following question: can epistemic uncertainty shape a transformer’s internal attention at inference time, without retraining or modifying pretrained weights?
We argue that effective uncertainty quantification requires uncertainty to actively shape attention, rather than merely annotate final predictions. For ambiguous inputs, a calibrated model should hedge internal reasoning instead of propagating early attention decisions deterministically. Existing stochastic inference methods expose uncertainty only at the output level and do not use it to modulate internal attention patterns (Fig.˜1C).
Motivated by this observation, we propose UAT-Lite, an inference-time framework for encoder-based pretrained transformer classifiers. UAT-Lite estimates token-level epistemic uncertainty via Monte Carlo sampling and uses it to downweight attention contributions from unstable tokens, without modifying pretrained weights or training objectives. Our main contributions are:
Uncertainty-Weighted Attention. An inference-time mechanism that modulates self-attention using token-level epistemic uncertainty estimated via Monte Carlo sampling.
Layer-Wise Uncertainty Attribution. A diagnostic variance decomposition that attributes predictive uncertainty across transformer depth.
Comprehensive Evaluation. Empirical evaluation across multiple NLP benchmarks demonstrating improved calibration, selective prediction, and robustness under distribution shift without degrading task performance.
UAT-Lite introduces no additional trainable parameters and operates entirely at inference time, requiring only a small number of stochastic forward passes. Post-hoc calibration methods such as temperature scaling are complementary rather than competing. The framework trades modest inference-time cost for improved reliability while avoiding the training and storage overhead of ensemble-based approaches. Our study focuses on encoder-based transformers; extending uncertainty-aware attention to decoder-only generative models is left for future work.
2 Preliminaries
Bayesian inference and uncertainty.
Bayesian analysis incorporates parameter uncertainty by averaging predictions over probable parameter values inferred from the training data. This perspective distinguishes epistemic uncertainty, which arises from limited data or insufficient coverage and reflects uncertainty about the model itself, from aleatoric uncertainty, which represents irreducible noise in the observations. In large neural models, epistemic uncertainty becomes particularly significant in the presence of semantic ambiguity or distributional shifts, where confident yet incorrect predictions frequently occur.
Monte Carlo (MC) Dropout.
We briefly recall MC Dropout as an approximate Bayesian inference technique used throughout this work. MC dropout retains dropout layers at inference time and performs multiple stochastic forward passes, yielding a predictive distribution whose variance estimates epistemic uncertainty and whose mean defines the final prediction. Conventional MC dropout treats uncertainty solely at the output level and does not influence internal representations or attention mechanisms, motivating methods that integrate uncertainty directly into the inference process.
Token-level uncertainty as a proxy signal.
While epistemic uncertainty is often measured at the output level, internal transformer computation is structured around token-level representations and attention interactions. Consequently, uncertainty estimates derived from stochastic embedding representations can serve as a lightweight proxy for identifying unstable or ambiguous token contributions. This perspective motivates the use of token-level uncertainty as a control signal for modulating attention during inference, without requiring architectural changes or retraining.
3 Related Work
Uncertainty estimation and calibration for neural networks have been widely studied, with post-hoc methods such as temperature scaling and Platt scaling (Guo et al., 2017; Platt et al., 1999) providing simple and effective confidence correction without altering model internals. Ensemble-based approaches, including Deep Ensembles (Lakshminarayanan et al., 2017), yield strong calibration but incur substantial training and storage overhead. Bayesian neural networks, stochastic inference methods, and recent Bayesian Transformer variants (Blundell et al., 2015; Gal and Ghahramani, 2016; Ritter et al., 2018; Fan et al., 2020; Zhang et al., 2021; Chen and Li, 2023) offer principled uncertainty modeling but typically require architectural modification, variational training, or task-specific retraining. In contrast, UAT-Lite integrates epistemic uncertainty directly into self-attention at inference time, preserving pretrained weights and standard training pipelines while enabling uncertainty-aware internal computation. A comprehensive discussion of related work and detailed comparisons is provided in App.˜G.
4 UAT-Lite
UAT-Lite is an uncertainty-aware extension of pretrained transformer classifiers designed to improve confidence calibration at inference time without modifying model architecture, training objectives, or learned parameters. The framework integrates approximate Bayesian uncertainty estimation into internal attention computation, allowing epistemic uncertainty to actively influence contextualization while preserving the base transformer’s parameterization. In addition, UAT-Lite introduces a diagnostic layer-wise variance decomposition that characterizes how predictive uncertainty evolves across transformer depth, enabling post-hoc analysis without affecting the forward pass. Final predictions are produced via stochastic inference with confidence-aware decision shaping. Overall, UAT-Lite preserves the training cost and parameter count of the underlying encoder, incurring only modest inference-time overhead.
4.1 Problem Setup and Notation
We consider an encoder-based pretrained transformer for classification. Given an input sequence , the model maps through stacked layers to contextual representations , which are converted to a predictive distribution by a task-specific classification head. In standard evaluation mode, pretrained transformers produce a single point estimate: dropout is disabled and the same input yields the same representations and prediction given fixed parameters. While effective for accuracy, this design provides no mechanism for epistemic uncertainty. Our goal is to introduce controlled inference-time stochasticity that allows uncertainty to influence internal attention without modifying training or parameters.
4.2 Estimating Uncertainty via MC Sampling
We estimate predictive uncertainty using MC dropout, a lightweight approximation to Bayesian inference. Dropout layers are retained during inference and stochastic forward passes are performed, inducing variability via random neuron masking. Aggregating these samples yields a predictive mean and an estimate of epistemic variability across stochastic passes during evaluation.
Uncertainty type.
Following Kendall and Gal (2017), predictive uncertainty decomposes into aleatoric and epistemic components. MC dropout primarily captures epistemic uncertainty through approximate posterior sampling during inference.
Component-specific dropout.
To balance representational stability and uncertainty estimation, we apply component-specific dropout during inference (0.1 for embeddings, 0.2 for attention layers, and 0.3 for feed-forward layers). These rates follow common practice in transformer fine-tuning and are kept fixed across tasks; calibration trends are stable under moderate variations of these values.
Single-pass sampling protocol.
We use a single-pass MC inference procedure in which the same stochastic forward passes are used both to (i) estimate token-level embedding uncertainty and (ii) obtain the final MC predictive distribution. As detailed in Algorithm 1, for each pass we sample dropout masks, compute embeddings , propagate the input through the encoder with uncertainty-weighted attention, and record logits for aggregation. Token-level uncertainty is computed by aggregating embedding variability across these passes and is reused consistently within the same inference run. This design avoids any separate embedding-only sampling stage and ensures that inference-time overhead scales linearly with .
4.3 Uncertainty-Weighted Self-Attention
We incorporate epistemic uncertainty directly into self-attention by modulating attention logits using token-level uncertainty estimates computed at the embedding layer during contextualization.
Token-level uncertainty estimation.
For each input token , we obtain stochastic embedding samples . Token-level uncertainty is defined as
where denotes the embedding dimensionality. We interpret as an embedding-level proxy for epistemic uncertainty induced by dropout.
This quantity is computed once per input and shared across all layers and attention heads. The computational cost is negligible relative to Monte Carlo methods, as it involves only per-token operations over the sampled embeddings and does not require additional forward passes.
Scale sensitivity.
The magnitude of remains well-behaved due to embedding normalization and averaging across dimensions. Sensitivity analysis over and (§˜D.1) shows stable calibration across a broad range of values; explicit normalization is therefore unnecessary.
Attention modulation.
Let , , and denote query, key, and value projections. Standard scaled dot-product attention logits are
We modulate these logits using token-level uncertainty:
where is a fixed uncertainty penalty. Attention probabilities are obtained via softmax:
Relation to prior work.
Design choices.
Uncertainty modulation is applied uniformly across layers and heads and affects only the key dimension in the attention mechanism. Unless otherwise stated, the uncertainty penalty (default ) is set once on development data and held fixed across tasks. §˜D.1 shows that calibration is stable across a broad range of and . The method introduces no additional trainable parameters beyond the base model.
4.4 Layer-Wise Uncertainty Attribution
We analyze how predictive uncertainty propagates across transformer depth under inference-time stochasticity. This analysis is diagnostic and does not affect forward computation or model behavior.
Layer-wise variance decomposition.
Using the law of total variance, total predictive variance can be decomposed into layer-specific contributions under inference-time stochasticity. A formal derivation is provided in App.˜A.
Theorem 1 (Layer-Wise Variance Decomposition).
Under inference-time stochasticity, total predictive variance can be expressed as the sum of layer-specific variance contributions.
Scope and interpretation.
This result follows from standard variance identities and is not a novel mathematical contribution. Its value lies in enabling diagnostic inspection of uncertainty evolution across transformer depth, which is unavailable in post-hoc calibration methods. Layer-wise contributions should be interpreted as approximate signals rather than strictly causal attributions.
4.5 Confidence-Aware Decision Shaping
MC dropout produces a distribution of stochastic predictions. The predictive mean is computed from the stochastic forward passes. Post-hoc temperature scaling is applied to the Monte Carlo mean logits, with the temperature parameter learned on the validation set to calibrate predictive confidence.
Selective prediction.
Uncertainty estimates enable abstention via confidence thresholds selected on the validation split of each dataset and applied uniformly at test time across all evaluated methods.
5 Experimental Setup
5.1 Datasets
We evaluate on five benchmarks spanning general NLP and clinical domains commonly used in prior calibration and robustness studies. All experiments focus exclusively on encoder-based pretrained transformer classifiers (BERT-family backbones), rather than decoder-only language models. General NLP datasets include SQuAD 2.0 Rajpurkar et al. (2018) (QA with unanswerable questions), MNLI Williams et al. (2018) (natural language inference with matched/mismatched splits for distribution shift analysis), and SST-2 Socher et al. (2013) (binary sentiment classification). Clinical benchmarks include MedQA Jin et al. (2021) and PubMedQA Jin et al. (2019), which are used to assess domain transfer rather than clinical validity.
5.2 Baselines, Metrics, and Training Overview
We evaluate our approach against BERT-base, MC Dropout, Temperature Scaling (TS), Deep Ensembles, and representative Bayesian last-layer methods. All models share the same pretrained backbone, fine-tuning recipe, and evaluation protocol; detailed baseline configurations are described in App.˜C.
Calibration is evaluated using Expected Calibration Error (ECE), alongside robustness under distribution shift and selective prediction metrics. Results are reported as mean standard deviation over five random seeds, with deterministic baselines exhibiting negligible variance. The proposed method introduces no additional trainable parameters and operates entirely at inference time.
Uncertainty penalty.
Unless otherwise stated, we use a single uncertainty penalty , set once on development data and held fixed across all tasks and datasets. §˜D.1 shows that calibration performance is stable across a broad range of and Monte Carlo budgets , indicating that the method does not rely on sensitive hyperparameter tuning in practice across the evaluated settings.
6 Results
6.1 General NLP Calibration Performance
Table 1 reports multi-seed calibration results (mean standard deviation over five seeds), computed using identical ECE binning and confidence definitions across methods. Deterministic baselines exhibit negligible variance, while stochastic methods show small but consistent variability.
UAT-Lite improves calibration relative to the unscaled fine-tuned BERT baseline, reducing average ECE from 0.117 to 0.094 (19.9% relative reduction). The largest gain is observed on MNLI (0.110 0.067), a task characterized by semantic ambiguity, where uncertainty-aware attention is particularly beneficial. These improvements persist relative to both uniform and component-specific MC Dropout, indicating that the gains arise from uncertainty-weighted attention rather than from stochasticity introduced by sampling.
Temperature scaling (TS) achieves the lowest in-domain ECE across all tasks, as expected for a post-hoc method explicitly optimized to minimize calibration error on held-out validation data. Our objective is complementary rather than competitive: while TS rescales output probabilities, it does not alter the internal representations or attention patterns that produce a prediction. In contrast, UAT-Lite introduces uncertainty-aware internal computation at inference time, allowing epistemic uncertainty to influence token interactions during contextualization. Because TS operates purely at the output level, it can be directly stacked on top of UAT-Lite without modification; we focus here on isolating the effects of uncertainty-aware attention and evaluate its benefits under distribution shift and selective prediction in subsequent sections.
Among Bayesian-style alternatives, SNGP achieves lower ECE but requires architectural modification and specialized training. Deep Ensembles provide a strong reference point (Avg ECE 0.089), but incur substantial training and storage overhead due to maintaining multiple model replicas. In contrast, UAT-Lite operates on a single pretrained transformer and introduces uncertainty handling entirely at inference time. As a result, it substantially narrows the calibration gap to Deep Ensembles while preserving deployment efficiency.
| Method | SQuAD (ECE ) | MNLI (ECE ) | SST-2 (ECE ) | Avg (ECE ) |
| BERT-base | 0.186 0.000 | 0.110 0.000 | 0.055 0.000 | 0.117 |
| MC Dropout (uniform ) | 0.173 0.001 | 0.090 0.001 | 0.039 0.001 | 0.101 |
| MC Dropout (component-specific) | 0.186 0.000 | 0.109 0.001 | 0.054 0.001 | 0.116 |
| Temp. Scaling | 0.057 0.000 | 0.028 0.000 | 0.029 0.000 | 0.038 |
| SNGP | 0.001 0.000 | 0.021 0.000 | 0.009 0.000 | 0.010 |
| VI-LastLayer | 0.119 0.001 | 0.096 0.000 | 0.136 0.002 | 0.117 |
| Deep Ensemble (5) | 0.168 0.000 | 0.060 0.000 | 0.037 0.000 | 0.089 |
| UAT-Lite | 0.163 0.002 | 0.067 0.001 | 0.051 0.004 | 0.094 |
| vs. Baseline | 12.5% | 39.0% | 6.6% | 19.9% |
6.2 Distribution Shift Robustness
We evaluate calibration robustness under distribution shift on MNLI by transferring models fine-tuned on the matched (in-domain) split to the mismatched (out-of-distribution) split. Results in Table 2 are reported as mean standard deviation over five seeds, with ECE computed consistently.
Absolute ECE values differ from the in-domain setting because MNLI shift evaluation is framed as a binary decision (entailment vs. non-entailment), and confidence is computed from the Monte Carlo–averaged predictive distribution using the maximum softmax probability. This affects the scale of ECE but preserves meaningful relative comparisons across methods.
BERT-base exhibits clear calibration drift under shift, with ECE increasing from the matched to the mismatched split (). In contrast, UAT-Lite yields lower ECE on both in-domain and out-of-domain splits and reduces calibration drift (), resulting in a substantially lower average ECE across the two conditions.
Importantly, this evaluation isolates the effect of uncertainty-aware internal computation under shift. Post-hoc calibration methods such as temperature scaling are intentionally omitted here, as they rescale output confidences without modifying the internal representations that give rise to distributional sensitivity. The improved stability observed for UAT-Lite therefore reflects the benefit of allowing epistemic uncertainty to modulate attention during contextualization, rather than relying solely on output-level correction. Overall, these results indicate that uncertainty-weighted attention produces more stable confidence estimates under matched-to-mismatched distribution shift, a setting where overconfident errors are particularly costly.
| Method | ID ECE | OOD ECE | ECE | Avg ECE |
| BERT-base | 0.2893 0.0000 | 0.3017 0.0000 | 0.0124 | 0.2955 |
| UAT-Lite | 0.0219 0.0007 | 0.0145 0.0007 | -0.0074 | 0.0182 |
6.3 Selective Prediction Performance
Table 3 reports selective prediction performance using AURC (lower is better), averaged over five seeds. All selection criteria are determined on validation data and held fixed across methods. Because the Baseline is deterministic, its variance across seeds is negligible, whereas UAT-Lite exhibits small but consistent variability across runs.
Selective prediction directly evaluates the quality of uncertainty estimates rather than calibration alone: a reliable model should abstain on difficult inputs while maintaining high coverage on confident predictions. Accordingly, this evaluation isolates whether uncertainty estimates meaningfully align with prediction risk under identical settings.
On MNLI, UAT-Lite achieves substantially higher selective coverage, increasing from 0.714 to 0.869, while reducing AURC from 0.164 to 0.053. This reflects a markedly improved risk–coverage trade-off, where abstention is applied preferentially to higher-risk inputs rather than uniformly.
On SQuAD 2.0, coverage increases from 0.501 to 0.690 and AURC decreases from 0.447 to 0.289, indicating more effective abstention behavior on ambiguous or unanswerable inputs. These gains demonstrate that uncertainty-weighted attention yields confidence estimates that better reflect underlying uncertainty.
Post-hoc calibration methods such as temperature scaling are intentionally omitted from this analysis, as they modify output probabilities without altering the confidence ranking that governs abstention behavior. These improvements reflect uncertainty-aware internal computation.
| Dataset | Method | Coverage@0.9 | Coverage@0.8 | Coverage@0.7 | AURC |
| MNLI | Baseline | 0.7139 0.0000 | 0.7527 0.0000 | 0.7926 0.0000 | 0.1640 0.0000 |
| MNLI | UAT-Lite | 0.8692 0.0003 | 0.9030 0.0006 | 0.9275 0.0010 | 0.0529 0.0002 |
| SQuAD 2.0 | Baseline | 0.5007 0.0000 | 0.5071 0.0000 | 0.5050 0.0000 | 0.4472 0.0000 |
| SQuAD 2.0 | UAT-Lite | 0.6904 0.0006 | 0.7047 0.0009 | 0.7163 0.0010 | 0.2892 0.0007 |
6.4 Computational Efficiency
Table 4 reports inference cost and calibration performance on SQuAD 2.0 under a single-accelerator, sequential inference setting. We focus on methods with clearly defined deployment behavior and directly measurable inference overhead. Predictive accuracy is not reported in this analysis, as all compared methods exhibit comparable task performance on SQuAD 2.0 and the purpose of this section is to isolate calibration behavior under differing inference-time costs.
UAT-Lite with incurs higher latency than the deterministic baseline (86.31 ms vs. 17.69 ms), reflecting the cost of repeated stochastic forward passes, and yields a throughput of 13.9 examples/sec. Under the same sequential evaluation protocol, this latency is comparable to that of a 5-model Deep Ensemble (88.80 ms), while requiring only a single pretrained model (109M parameters) rather than storing and maintaining multiple independently trained replicas (547M parameters).
Temperature scaling adds negligible runtime overhead relative to the baseline, but operates purely as an output-level post-hoc calibration method and does not alter internal computation. Overall, UAT-Lite incurs additional inference-time cost in exchange for improved calibration, while avoiding the substantial training and storage overhead associated with ensemble-based approaches and preserving a single-model deployment footprint effectively.
| Method | Params | Samples | Latency (ms) | Memory (GB) | Throughput | ECE |
| BERT-base | 109M | 1 | 17.69 | 1.78 | 68.9 | 0.186 |
| Temperature Scaling | 109M | 1 | 17.71 | 1.78 | 69.0 | 0.057 |
| Deep Ensemble (5) | 547M | 5 | 88.80 | — | — | 0.168 |
| UAT-Lite | 109M | 10 | 86.31 | 1.81 | 13.9 | 0.163 |
6.5 Negation-Aware Attention
Fig.˜3 qualitatively illustrates how uncertainty-weighted attention modulates token interactions under linguistic negation. We compare attention patterns from a standard transformer and UAT-Lite under negated linguistic constructions.
While absolute attention maps remain similar due to softmax normalization, the difference heatmap reveals small but structured redistributions. Uncertainty-weighted attention slightly downweights sentiment-bearing tokens (e.g., “positively”) while reallocating attention toward negation cues and predicate structure associated with the phrase “did not respond” in context.
These changes are intentionally modest and should be interpreted as directional rather than large-scale attention reconfiguration. Accordingly, this example is illustrative rather than a precise token-level attribution. Corpus-level analysis shows that similar directional trends recur across negated and uncertainty-sensitive sentences (see §˜D.4 for additional qualitative evidence).
6.6 Ablation Study
Fig.˜4 decomposes UAT-Lite into its constituent components to isolate their individual contributions. Uncertainty-weighted attention alone reduces ECE from 0.169 to 0.134 (21.1% relative reduction), indicating that attention modulation based on epistemic uncertainty is the primary driver of calibration improvement under controlled comparisons. Embedding-level uncertainty alone slightly degrades calibration, suggesting that stochasticity without uncertainty-aware computation is insufficient. In contrast, combining uncertainty-weighted attention with confidence-aware decision shaping yields the largest improvement (28.2% relative reduction), exceeding either component in isolation and demonstrating their complementarity.
6.7 Generalization Across BERT Models
We evaluate generalization across BERT-based encoder architectures spanning general-purpose, biomedical, clinical, and scientific domains (App.˜E). Under matched evaluation protocols, UAT-Lite improves calibration across these pretrained backbones, yielding an average relative ECE reduction of approximately 32%.
The largest gains occur for mid-sized and domain-specialized models, including BERT-base on SQuAD, BioBERT on PubMedQA, and ClinicalBERT on MedQA, where epistemic uncertainty arising from semantic ambiguity is prominent. Improvements are smaller for models with strong baseline calibration (e.g., RoBERTa-base on MNLI) and diminish or reverse for very small or highly overparameterized models.
These results indicate that uncertainty-aware attention generalizes reliably within the BERT encoder family. Extending the mechanism to fundamentally different architectures, such as decoder-only or state-space models, remains an important direction for future work.
Additional analyses.
Additional analyses on adversarial robustness, linguistic uncertainty, clinical-domain transfer, and layer-wise uncertainty diagnostics are provided in the App.˜D.
7 Conclusion
We introduced UAT-Lite, an inference-time framework that integrates epistemic uncertainty into transformer computation through uncertainty-weighted attention and a layer-wise variance diagnostic. Across NLP benchmarks, UAT-Lite reduces Expected Calibration Error relative to unscaled fine-tuned BERT (e.g., average ECE 0.117 0.094) and improves selective prediction behavior, including higher coverage at fixed confidence thresholds and lower AURC. The method also exhibits stable calibration under distribution shift in MNLI matchedmismatched transfer. While inference-time Monte Carlo sampling incurs additional latency, the approach avoids the training and storage costs of ensemble-based alternatives and requires no modification of pretrained weights. Overall, UAT-Lite provides a complementary mechanism for uncertainty-aware transformer inference, bridging the gap between post-hoc calibration and Bayesian modeling for transformers.
Limitations
Our framework incurs additional computational cost due to Monte Carlo sampling, with a latency increase at . This overhead is appropriate for offline, batch, or document-level applications where confidence reliability is prioritized but is prohibitive for latency-critical settings. In practical deployments, this trade-off suggests that uncertainty-aware inference should be selectively enabled for high-risk or ambiguous inputs, rather than applied uniformly across all predictions. Accordingly, UAT-Lite is best viewed as an intermediate solution between lightweight post-hoc calibration and more expensive ensemble-based approaches, rather than a universal replacement.
Finally, on clinical benchmarks (MedQA and PubMedQA), calibration gains are modest, reflecting dataset properties such as factual emphasis and limited ambiguity; these experiments are therefore treated as domain-transfer stress tests rather than clinical validation.
Appendix
Appendix A Mathematical Details
In this section, we provide the formal mathematical foundations underlying the uncertainty modeling and diagnostic analyses used in UAT-Lite. We first introduce a layer-wise variance decomposition that attributes predictive uncertainty across transformer depth under stochastic inference. We then provide a proof of the decomposition, clarify its assumptions, and relate Monte Carlo dropout to approximate Bayesian inference. Finally, we discuss practical and interpretive implications of layer-wise uncertainty attribution for calibration, intervention, and uncertainty-aware inference. These analyses are intended as diagnostic and explanatory tools that complement the empirical results in the main paper, rather than as claims of exact posterior inference.
A.1 Layer-Wise Uncertainty Decomposition
To better understand how predictive uncertainty emerges within the transformer, we analyze its distribution across layers under stochastic inference.
Normalized variance contributions.
To facilitate interpretability and ensure internal consistency, we report normalized layer-wise variance contributions, defined as
such that contributions across layers sum to . All percentages reported below correspond to these normalized values.
Table 5 presents representative layer-wise uncertainty decompositions for answerable and unanswerable SQuAD examples, illustrating how predictive uncertainty is distributed across transformer depth rather than manifesting solely at the output layer. We emphasize that these examples are illustrative diagnostics rather than statistical summaries, and are intended to provide qualitative insight into uncertainty propagation rather than population-level guarantees.
For the unanswerable example, uncertainty grows progressively through the self-attention stack. While embedding-level contributions remain modest (3.6%), uncertainty increases substantially in mid and late layers, with the largest contribution arising from layers 9–11 (12.5%). This pattern is consistent with uncertainty emerging during later reasoning stages when sufficient evidence is unavailable, rather than arising only at the final decision layer.
In contrast, the answerable example exhibits a similar normalized uncertainty distribution but with a different internal structure. Mid-layer attention contributes more strongly to evidence aggregation (8.7%), while late-layer uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
These cases illustrate that aggregate predictive variance alone is insufficient to characterize model reliability. Instead, the distribution of uncertainty across layers provides insight into whether uncertainty arises from ambiguous inputs or from higher-level reasoning processes.
| Layer | Component | Normalized Var. | % | Interpretation |
| SQuAD Unanswerable | ||||
| Layer 0 | Embedding | 0.036 | 3.6% | Word ambiguity |
| Layer 1–4 | Attention (early) | 0.054 | 5.4% | Context formation |
| Layer 5–8 | Attention (mid) | 0.088 | 8.8% | Escalating uncertainty |
| Layer 9–11 | Attention (late) | 0.125 | 12.5% | Uncertainty consolidation |
| Layer 12 | Output | 0.021 | 2.1% | Final decision |
| Total | 1.000 | 100% | Correctly uncertain | |
| SQuAD Answerable | ||||
| Layer 0 | Embedding | 0.037 | 3.7% | Clear input |
| Layer 1–4 | Attention (early) | 0.054 | 5.4% | Evidence aggregation |
| Layer 5–8 | Attention (mid) | 0.087 | 8.7% | Confidence reinforcement |
| Layer 9–11 | Attention (late) | 0.126 | 12.6% | Confident reasoning |
| Layer 12 | Output | 0.023 | 2.3% | Final decision |
| Total | 1.000 | 100% | Confident prediction | |
A.2 Proof of Theorem 1
Proof.
We derive the layer-wise variance decomposition by applying the law of total variance recursively across transformer layers.
Setup and notation.
Let denote the model’s final prediction (e.g., class probability or logit), and let denote the hidden representation at layer corresponding to the [CLS] token. Stochasticity is introduced via Monte Carlo dropout at inference time.
Each transformer layer implements a stochastic mapping
where denotes dropout-induced noise at layer . We assume finite second moments for all noise distributions but do not assume Gaussianity or independence across layers.
Layer-wise variance contribution.
We define the variance contribution of layer as
Recursive variance decomposition.
Applying the law of total variance at the final layer yields
Applying the same decomposition recursively yields an additive expansion over layers:
Interpretation.
This decomposition provides an attribution of predictive uncertainty across layers under conditional expectations. While inter-layer dependencies may introduce higher-order interaction effects, the decomposition remains useful as an interpretable diagnostic of where uncertainty is amplified during inference. ∎
A.3 Bayesian Interpretation of MC Dropout
Monte Carlo dropout provides a computationally efficient approximation to Bayesian inference in deep neural networks Gal and Ghahramani (2016). Rather than maintaining a full posterior distribution over model parameters, dropout induces stochasticity that can be interpreted as sampling from an implicit variational posterior under standard assumptions.
Bayesian predictive distribution.
In Bayesian inference, predictions marginalize over uncertainty in model parameters:
Monte Carlo approximation via dropout.
MC dropout approximates this marginalization by retaining dropout during inference and performing stochastic forward passes:
Uncertainty interpretation.
The variability across predictions reflects epistemic uncertainty arising from model and data limitations, which is reducible with additional information.
Relevance to UAT-Lite.
In UAT-Lite, these uncertainty estimates are explicitly propagated through attention mechanisms, enabling uncertainty-aware inference without modifying training objectives or learned parameters.
A.4 Implications of Layer-Wise Uncertainty
Attribution.
Decomposing predictive variance across transformer depth enables identification of where uncertainty is amplified during inference. As shown in Table 5, unanswerable SQuAD examples concentrate substantial uncertainty in late layers (9–11), consistent with ambiguity emerging during final reasoning stages. Answerable examples exhibit comparable late-layer contributions, but with uncertainty reflecting confident task-level reasoning rather than unresolved ambiguity.
Intervention.
Layer-wise attribution suggests that calibration strategies could in principle target high-variance components. Our ablation study (Fig.˜4) shows that uncertainty introduced at the attention level yields a 21.1% improvement in calibration, while embedding-level uncertainty contributes negatively (-3.1%), supporting selective intervention rather than uniform stochasticity.
Interpretability.
Unlike post-hoc calibration methods that produce a single uncertainty score, this decomposition reveals how uncertainty evolves during reasoning. For complex syntactic constructions, uncertainty is elevated in early and mid layers due to structural processing, while late layers consolidate the final decision.
A.5 Practical Inference-Time Procedure
At test time, Monte Carlo dropout produces multiple stochastic forward passes of the same pretrained transformer. These samples yield a predictive mean and variance. The mean is used as the final prediction, while the variance quantifies epistemic uncertainty.
Uncertainty estimates influence inference by modulating attention behavior under representation ambiguity and enabling confidence-aware decision shaping at the output level. All uncertainty mechanisms operate strictly at inference time and do not modify training dynamics or learned parameters.
Appendix B Inference-Time Algorithm
Inference-time algorithm.
Algorithm 1 summarizes the inference-time procedure used by UAT-Lite. Uncertainty estimation and prediction are performed using the same stochastic forward passes, without any separate embedding-only sampling stage.
Appendix C Baselines and Metrics
We compare against BERT-base (110M parameters), MC Dropout (, ), Temperature Scaling (TS), and Deep Ensembles (5 independently fine-tuned models). Temperature scaling is treated as a strong post-hoc calibration baseline that operates purely at the output level and does not modify internal model computation.
Because TS rescales logits without affecting representations, it is compatible with uncertainty-aware inference and could in principle be applied on top of UAT-Lite. In this work, however, we report TS as a standalone baseline and do not include a stacked (UAT-Lite + TS) variant in the main tables, in order to isolate the effects of uncertainty-aware internal computation from output-level rescaling.
All baselines use the same pretrained backbone, fine-tuning recipe, and evaluation code unless otherwise stated.
Bayesian baselines beyond inference-time methods.
Methods such as SWAG and Laplace approximations estimate parameter-space uncertainty through task-specific posterior fitting or fine-tuning (Talman et al., 2023; Ritter et al., 2018). Because our scope is inference-time calibration of pretrained transformers without additional training or posterior fitting, we do not evaluate these approaches as direct baselines. We discuss them in Related Work (App.˜G) as complementary directions.
Expected Calibration Error (ECE).
Calibration is measured using Expected Calibration Error (ECE):
where predictions are partitioned into fixed-width confidence bins over . For each bin, denotes empirical accuracy and denotes the mean predicted confidence. Here denotes the number of confidence bins used for calibration evaluation, distinct from the number of Monte Carlo samples used for uncertainty estimation. Lower ECE indicates better calibration.
Confidence definition.
For classification tasks (MNLI, SST-2, MedQA), confidence is defined as the maximum softmax probability of the predicted class, computed after aggregating Monte Carlo samples. For SQuAD 2.0, confidence corresponds to the predicted answerability probability (unanswerable vs. answerable), following standard practice in uncertainty-aware question answering.
SQuAD 2.0 evaluation protocol.
We evaluate SQuAD 2.0 on the answerability prediction task, where the model must determine whether a question is answerable given the context. This binary classification task is well-suited for calibration analysis. We report answerability prediction accuracy alongside ECE in Table 1 . Note that span-level extraction (predicting answer start/end positions) is a separate task and orthogonal to our uncertainty calibration contribution. Our focus is on calibrated confidence scores for the answerability decision, not optimizing span extraction performance.
Distribution shift robustness metrics.
To assess calibration stability under distributional shift, we evaluate models on both in-domain (ID) and out-of-distribution (OOD) data and report two complementary metrics. First, we measure calibration drift:
where smaller (or negative) values indicate improved robustness under shift. Second, we report a calibration robustness score defined as the average ECE across ID and OOD settings:
with lower values indicating more stable calibration across domains.
Selective prediction metrics.
We additionally report selective coverage at fixed confidence thresholds and the Area Under the Risk–Coverage curve (AURC; lower is better). Confidence thresholds are selected on the validation set and held fixed at test time for all methods.
MC Dropout baseline configuration.
The MC Dropout baseline uses a uniform dropout rate of across all layers, consistent with prior calibration studies. While more aggressive or component-specific dropout tuning could further improve MC Dropout performance, such tuning increases implementation complexity and reduces comparability with standard baselines. In contrast, UAT-Lite explicitly leverages structured uncertainty within the attention mechanism rather than relying solely on increased stochasticity. This comparison isolates the effect of uncertainty-aware attention rather than raw dropout intensity.
Component-specific MC Dropout control.
To explicitly disentangle the effect of structured stochasticity from uncertainty-aware attention, we additionally evaluate an MC Dropout baseline using the same component-specific dropout rates as UAT-Lite (0.1 embeddings / 0.2 attention / 0.3 feed-forward), but without attention modulation. As shown in Table 1, this control does not improve calibration relative to uniform MC Dropout or the BERT baseline, indicating that the gains of UAT-Lite arise from uncertainty-weighted attention rather than tuned dropout alone.
Reporting conventions and fairness.
Within each table, all methods are evaluated using the same data splits, evaluation code, confidence definitions, and Monte Carlo budget (where applicable). ECE is computed with fixed-width bins for all methods. Temperature scaling (and UAT-Lite +TS) fits a single scalar temperature on the validation set only, and is applied at test time without any access to test labels. OOD calibration metrics (e.g., ECE and Robustness) are computed using the identical pipeline on ID and OOD sets to avoid binning or confidence-definition mismatch.
Random seeds and variability.
Main-paper results (general calibration, distribution-shift robustness, and selective prediction) are reported as mean standard deviation over five random seeds. Appendix diagnostic analyses—including qualitative attention visualizations, representative layer-wise uncertainty decompositions, and linguistic or adversarial case studies—are reported using a single seed to reduce computational cost and are explicitly labeled as single-seed in the corresponding captions. Deterministic baselines exhibit negligible variance because inference is fully deterministic when evaluated from a fixed fine-tuned checkpoint reused across seeds.
Appendix D Additional Experimental Results
This section presents additional analyses that complement the main experimental results by providing deeper insight into the behavior of UAT-Lite in specialized and fine-grained settings. Specifically, we examine calibration performance in clinical benchmarks, where decision reliability is critical, and conduct a detailed linguistic analysis to assess how uncertainty-aware attention responds to diverse sources of ambiguity, structure, and semantic inconsistency. Together, these experiments offer a more nuanced understanding of the strengths and limitations of the proposed framework beyond aggregate calibration metrics.
D.1 Sensitivity Analysis of and
We examine the sensitivity of UAT-Lite to its two primary inference-time hyperparameters: the uncertainty weighting coefficient and the number of Monte Carlo samples . A small, representative grid is evaluated with and on SQuAD 2.0, which is particularly sensitive to epistemic uncertainty due to the presence of both answerable and unanswerable questions. Similar stability trends were observed on MNLI and SST-2 and are omitted for brevity. We report ECE and predictive accuracy to assess both calibration quality and task performance. Accuracy is reported here solely to verify that calibration gains do not arise from performance degradation.
Across all configurations, calibration performance remains stable, with ECE values confined to a narrow interval between 0.064 and 0.070 (mean , standard deviation ). No sharp sensitivity to either or is observed. Increasing the number of Monte Carlo samples yields only marginal calibration improvements beyond , indicating diminishing returns from additional stochastic passes.
Predictive accuracy remains effectively unchanged across the grid, varying by less than 0.5% absolute, suggesting that uncertainty-weighted attention primarily affects calibration rather than accuracy. Overall, these results indicate that UAT-Lite does not rely on precise hyperparameter tuning to achieve effective calibration and is robust across a broad range of reasonable inference-time settings. This stability supports the practicality of the method in deployment scenarios where extensive hyperparameter search is undesirable.
| ECE | Accuracy | ||
| 0.1 | 3 | 0.0699 | 0.7075 |
| 0.1 | 5 | 0.0676 | 0.7075 |
| 0.1 | 10 | 0.0649 | 0.7109 |
| 0.5 | 3 | 0.0703 | 0.7059 |
| 0.5 | 5 | 0.0657 | 0.7096 |
| 0.5 | 10 | 0.0682 | 0.7068 |
| 1.0 | 3 | 0.0669 | 0.7090 |
| 1.0 | 5 | 0.0655 | 0.7102 |
| 1.0 | 10 | 0.0640 | 0.7110 |
Summary. ECE mean , std ; min , max (range ).
D.2 Clinical Domain Calibration
Table 7 reports calibration and accuracy on two clinical benchmarks with distinct task structures and uncertainty profiles: MedQA and PubMedQA. We emphasize that these clinical datasets are used only to assess domain transfer of uncertainty-aware attention to medically specialized language distributions, and should not be interpreted as evidence of clinical decision reliability. In particular, what constitutes “good uncertainty” in clinical settings is often decision-dependent and may involve asymmetric error costs (e.g., prioritizing recall for critical conditions), which are not explicitly modeled in our evaluation.
On MedQA, Deep Ensembles achieve the lowest ECE (0.007) and the highest accuracy (0.268), reflecting the benefit of model diversity in discrete, fact-centric multiple-choice medical question answering. In contrast, UAT-Lite exhibits a slightly higher ECE (0.040) with marginal accuracy improvement over the baseline. This behavior is expected, as MedQA primarily evaluates factual recall with limited contextual ambiguity, constraining the benefits of uncertainty-aware attention. Importantly, UAT-Lite achieves competitive performance without the substantial computational and storage overhead associated with training and maintaining multiple model replicas.
A different trend emerges on PubMedQA. While Deep Ensembles substantially reduce ECE (0.051), this improvement coincides with a marked drop in accuracy (0.520), indicating overly conservative confidence estimates. In contrast, UAT-Lite attains the highest accuracy (0.640) while maintaining calibration comparable to the baseline. This suggests that integrating uncertainty directly into the attention mechanism can improve evidence aggregation in long, semantically dense biomedical abstracts, where uncertainty arises from contextual interpretation rather than discrete fact retrieval.
Overall, these results highlight a trade-off between ensemble-based calibration and inference-time uncertainty integration. While ensembles excel in fact-centric clinical QA, UAT-Lite offers a more favorable accuracy–calibration balance on tasks characterized by semantic ambiguity and long-context reasoning, without incurring additional training or parameter costs. Accordingly, we treat the clinical-domain results as a domain-transfer generalization stress test rather than a primary claim about clinical deployment.
| Dataset | Method | ECE | Accuracy |
| MedQA | Baseline | 0.033 | 0.246 |
| Temperature Scaling | 0.023 | 0.246 | |
| Deep Ensemble (5) | 0.007 | 0.268 | |
| UAT-Lite | 0.040 | 0.249 | |
| PubMedQA | Baseline | 0.114 | 0.620 |
| Temperature Scaling | 0.128 | 0.620 | |
| Deep Ensemble (5) | 0.051 | 0.520 | |
| UAT-Lite | 0.103 | 0.640 |
D.3 Character-Level Adversarial Robustness
Table 8 evaluates robustness under character-level perturbations on SST-2. We apply random 5% character swaps to simulate typographical errors and adversarial perturbations.
Both baseline BERT and UAT-Lite exhibit substantial accuracy degradation under attack, with baseline accuracy dropping from 0.920 to 0.816 and UAT-Lite from 0.922 to 0.812. Calibration also degrades substantially, with ECE increasing from 0.062 to 0.133 for baseline and from 0.055 to 0.132 for UAT-Lite.
While UAT-Lite maintains slightly better clean accuracy (0.922 vs. 0.920) and clean calibration (ECE 0.055 vs. 0.062), the robustness under attack is comparable between methods. This indicates that uncertainty-aware attention improves calibration on clean data but provides limited additional protection against adversarial perturbations.
These results are consistent with prior work showing that calibration methods address epistemic uncertainty but do not directly defend against adversarial attacks Goodfellow et al. (2014). Future work should explore combining UAT-Lite with specialized adversarial defenses such as adversarial training or certified robustness techniques.
| Attack | Method | Clean Acc | Attack Acc | Clean ECE | Attack ECE |
| Character Swap | Baseline | 0.920 | 0.816 | 0.062 | 0.133 |
| Character Swap | UAT-Lite | 0.922 | 0.812 | 0.055 | 0.132 |
D.4 Full Linguistic Uncertainty Analysis
Table 9 provides a fine-grained linguistic analysis of how UAT-Lite responds to diverse sources of uncertainty, spanning lexical, syntactic, epistemic, and logical phenomena.
For negation, UAT-Lite remains uniformly confident across affirmative, negated, and clinical negation cases, yielding zero estimated uncertainty. This behavior indicates negation invariance, which is desirable for explicit clinical negation (e.g., “not stable”) but also suggests limited sensitivity to polarity reversal in non-clinical contexts. This invariance is consistent with the attention analysis in Fig.˜3, which shows that uncertainty-aware attention redistributes weight conservatively toward negation and predicate structure without inducing elevated output-level uncertainty for clear clinical negation.
In cases of lexical ambiguity, the framework exhibits mixed behavior. When sufficient contextual disambiguation is present (“bank account”), UAT-Lite appropriately increases uncertainty. However, for genuinely ambiguous expressions (“bank closed”), the model remains highly confident, indicating overconfidence under unresolved polysemy. Partial sensitivity is observed for polysemous terms such as “bat,” where uncertainty increases but does not fully reflect ambiguity.
A similar pattern arises for epistemic hedges. Strong assertions (“definitely”) and weak hedges (“probably,” “possibly”) are treated with near-maximal confidence, suggesting hedge insensitivity. In contrast, explicit modal uncertainty (“might”) leads to a measurable increase in uncertainty, indicating that the model responds primarily to strong lexical uncertainty markers rather than gradated epistemic cues.
For syntactic complexity, UAT-Lite demonstrates heightened sensitivity. Medium-depth syntactic structures produce the largest uncertainty increase, reflecting increased structural processing difficulty. Interestingly, deeper structures elicit a non-monotonic response, suggesting that the model may partially adapt to highly regular but complex constructions rather than uniformly increasing uncertainty with depth.
Finally, under logical contradictions, UAT-Lite appropriately increases uncertainty for inconsistent inputs (“healthy dying”) while maintaining high confidence for coherent statements. This indicates effective detection of semantic inconsistency, consistent with the framework’s uncertainty propagation across layers.
| Phenomenon | Trigger | Base Conf | MB Conf | Unc | Diagnostic |
| Negation | |||||
| Positive | excellent | 0.53 | 1.00 | 0.00 | Negation-invariant |
| Negated | not excellent | 0.52 | 1.00 | 0.00 | Negation-invariant |
| Clinical | not stable | 0.55 | 1.00 | 0.00 | Confident clinical negation |
| Lexical Ambiguity | |||||
| Clear context | bank account | 0.51 | 0.70 | 0.30 | Appropriate ambiguity |
| Ambiguous | bank closed | 0.55 | 0.99 | 0.01 | Overconfident ambiguity |
| Polysemous | bat (animal/tool) | 0.54 | 0.83 | 0.17 | Partial polysemy sensitivity |
| Epistemic Hedges | |||||
| Definite | definitely | 0.51 | 1.00 | 0.00 | Strong assertion |
| Probable | probably | 0.53 | 0.99 | 0.01 | Hedge-insensitive |
| Possible | possibly | 0.52 | 0.99 | 0.01 | Hedge-insensitive |
| Uncertain | might | 0.51 | 0.85 | 0.15 | Uncertainty-aware |
| Syntactic Complexity | |||||
| Simple | depth=2 | 0.55 | 0.82 | 0.18 | Low structural load |
| Medium | depth=5 | 0.53 | 0.54 | 0.46 | Complexity-sensitive |
| Complex | depth=8 | 0.53 | 0.83 | 0.17 | Non-monotonic response |
| Contradictions | |||||
| Consistent | healthy, stable | 0.52 | 0.99 | 0.01 | Coherent input |
| Contradictory | healthy dying | 0.53 | 0.90 | 0.10 | Contradiction-aware |
Appendix E Cross-Model & Scaling Analysis
Scope clarification. All models evaluated in this section are encoder-only, BERT-derived transformer architectures. Generalization to fundamentally different architectures (e.g., decoder-only models such as GPT or LLaMA, or state-space models) is outside the scope of the current study and is left for future work.
E.1 Generalization Across BERT Backbones
Table 10 evaluates whether the calibration benefits of UAT-Lite generalize across BERT-based transformer backbones pretrained on heterogeneous corpora and applied to domain-specific benchmarks. The evaluated models span general-purpose (BERT-base, RoBERTa-base), clinical (ClinicalBERT), biomedical (BioBERT), and scientific (SciBERT) representations, thereby testing robustness across substantially different linguistic and semantic distributions.
Across evaluated domains, UAT-Lite reduces ECE for most pretrained backbones, yielding an average relative improvement of approximately 32% (Table 10).
The strongest gain is observed for BERT-base on SQuAD, where ECE decreases by 52% (0.130 0.062). This result is particularly significant because SQuAD contains both answerable and unanswerable questions, making it highly sensitive to epistemic uncertainty. The large improvement indicates that uncertainty-weighted attention enables the model to better recognize ambiguity during evidence aggregation rather than producing overconfident predictions.
In domain-specialized settings, BioBERT on PubMedQA achieves a 40% reduction in ECE, suggesting that uncertainty-aware attention is especially beneficial for long, evidence-dense biomedical text that requires integrating information across multiple sentences. ClinicalBERT on MedQA shows a 25% improvement, indicating that the method remains effective even for fact-oriented medical question answering, although gains are smaller due to relatively strong baseline calibration.
By contrast, RoBERTa-base on MNLI exhibits only a marginal 2% improvement. MNLI is a large, high-resource dataset with strong pretrained representations and comparatively well-calibrated baselines, leaving limited room for additional calibration gains. SciBERT on SQuAD similarly shows a modest 10% reduction, suggesting that pretraining on highly specialized scientific corpora may already encode task-relevant uncertainty patterns.
| Base Model | Domain | Dataset | Baseline ECE | +UAT-Lite ECE | Improvement |
| BERT-base | General | SQuAD | 0.130 | 0.062 | 52% |
| RoBERTa-base | General | MNLI | 0.040 | 0.039 | 2% |
| ClinicalBERT | Medical | MedQA | 0.049 | 0.037 | 25% |
| BioBERT | Biomedical | PubMedQA | 0.120 | 0.072 | 40% |
| SciBERT | Scientific | SQuAD | 0.092 | 0.083 | 10% |
| Average | — | — | 0.086 | 0.059 | 32% |
E.2 Effect of Model Capacity on Calibration
Table 11 analyzes the effect of model capacity on calibration improvement by applying UAT-Lite across different sizes of the BERT architecture, ranging from BERT-tiny (4.4M parameters) to BERT-large (335.1M parameters). Unlike Table 10, which varies domains and pretraining corpora, this analysis holds the architecture family fixed and isolates the role of parameter scale.
Small and mid-sized models benefit substantially from uncertainty-weighted attention. BERT-mini achieves the largest relative improvement, with a 58% reduction in ECE, followed by BERT-base with an approximately 53% improvement (52–53% depending on rounding and binning resolution). These models exhibit sufficient representational capacity to express epistemic uncertainty, while still relying on attention mechanisms that benefit from explicit uncertainty modulation.
BERT-tiny and BERT-small show more modest improvements (12% and 15%, respectively). In very small models, limited representational capacity constrains the extent to which uncertainty-aware reasoning can reshape internal attention patterns, resulting in smaller but still consistent calibration gains.
In contrast, BERT-large exhibits degraded calibration, with ECE increasing by 13% when UAT-Lite is applied. We hypothesize that in highly overparameterized models, stochastic perturbations introduced via Monte Carlo sampling may disrupt already well-formed attention patterns, leading to excessive variance propagation rather than meaningful uncertainty modeling. This observation suggests that uncertainty-weighted attention is most effective in the regime of small to mid-sized transformers, where model capacity and stochasticity appear to be better balanced.
Latency measurements further contextualize this behavior. While inference-time overhead remains moderate for mid-sized models (e.g., 7 ms 39 ms for BERT-base), it increases substantially for larger architectures (26 ms 134 ms for BERT-large), reinforcing the trade-off between calibration gains and computational cost.
| Model | Params | Baseline ECE | UAT-Lite ECE | Rel. | Latency |
| BERT-tiny | 4.4M | 0.070 | 0.061 | 12% | 1 5 |
| BERT-mini | 11.2M | 0.131 | 0.056 | 58% | 1 8 |
| BERT-small | 28.8M | 0.095 | 0.081 | 15% | 1 9 |
| BERT-base | 109.5M | 0.139 | 0.066 | 53% | 7 39 |
| BERT-large | 335.1M | 0.152 | 0.172 | 13% | 26 134 |
Appendix F Implementation and Reproducibility
This section characterizes the computational and implementation properties of UAT-Lite. We first analyze its asymptotic and practical computational complexity relative to standard transformer inference and ensemble-based uncertainty methods. We then quantify inference-time overhead introduced by Monte Carlo sampling and describe the experimental setup, hardware configuration, and reproducibility details required to replicate all reported results. Throughout, we emphasize that UAT-Lite operates strictly at inference time and does not modify pretrained parameters or training cost.
F.1 Computational Complexity Analysis
UAT-Lite preserves the parameter count and training cost of the base transformer. All uncertainty-aware mechanisms operate strictly at inference time and do not introduce additional trainable parameters or modify learned weights.
Additional computational cost arises solely from Monte Carlo (MC) sampling during inference:
Parameters: identical to the base model (e.g., 109M for BERT-base)
Training cost: unchanged
Calibration cost: one scalar temperature per task (optional)
Inference cost: stochastic forward passes (typically )
From a complexity perspective, inference-time cost scales linearly with the number of Monte Carlo samples . In contrast to ensemble-based approaches, which require training, storing, and evaluating multiple independent models, UAT-Lite incurs no additional training or memory overhead. As a result, its total computational footprint remains substantially lower than Deep Ensembles for comparable calibration gains.
F.2 Inference-Time Overhead
While the previous subsection characterizes asymptotic complexity, we now report empirical inference-time overhead under realistic deployment conditions.
At test time, UAT-Lite performs stochastic forward passes through the same pretrained model. When executed sequentially, runtime increases approximately linearly with . For BERT-base, this corresponds to a measured slowdown of approximately relative to deterministic inference when using .
Despite this overhead, inference-time cost remains practical for many deployment settings, particularly in scenarios where reliability and calibrated confidence are prioritized over raw throughput. Importantly, this cost is incurred only when uncertainty-aware inference is enabled and can be adjusted dynamically by reducing for high-confidence inputs, as discussed in Section Limitations.
F.3 Reproducible Experimental Setup
This subsection summarizes the experimental setup and implementation details necessary to reproduce our results. All experiments follow a consistent architecture, training configuration, and evaluation protocol across datasets.
Model architecture.
Base model: BERT-base-uncased
Number of parameters: 110M
Transformer layers: 12
Hidden dimension: 768
Attention heads: 12
Training hyperparameters.
Optimizer: AdamW with standard settings
Learning rate:
Learning rate schedule: linear warmup for first 500 steps, followed by linear decay
Batch size: 16
Number of epochs: 3
Weight decay: 0.01
Gradient clipping: 1.0
UAT-Lite configuration.
Inference via Monte Carlo sampling with stochastic forward passes
Component-specific dropout rates:
Embedding layer: 0.1
Attention layers: 0.2
Feed-forward layers: 0.3
Uncertainty weighting coefficient: (set once on development data and held fixed across tasks)
Sensitivity analysis for and reported in §˜D.1
F.4 Hardware, Latency, and Code Availability
Compute resources and latency measurement.
All experiments are conducted on a single NVIDIA A100 GPU (40GB). Unless otherwise stated, latency is measured as end-to-end forward-pass wall-clock time per example, averaged over 200 runs after 50 warm-up iterations. We use fixed sequence lengths (128 for SST-2 and MNLI; task-standard truncation for SQuAD answerability) and batch size 1.
For stochastic methods, latency is reported under the same Monte Carlo budget used during evaluation. Reported slowdowns reflect the ratio between stochastic inference and deterministic baseline inference under identical conditions.
Code availability.
To ensure reproducibility during the anonymous review process, we provide an anonymized repository containing the complete implementation, including training scripts, evaluation pipelines, and configuration files. All experiments are implemented using HuggingFace Transformers, PyTorch, and Python. The repository is available at: https://github.com/eliashossain001/uq_decomposition/tree/main.
Appendix G Supplementary Related Work
This section provides an extended discussion of prior work on calibration, uncertainty estimation, and uncertainty-aware Transformer modeling. Because the main paper focuses on method design and empirical evaluation rather than a comprehensive survey, we defer detailed comparisons to this supplementary section. We organize related work into three themes: (i) post-hoc calibration and ensemble-based uncertainty estimation, (ii) Bayesian Transformers and stochastic inference methods, and (iii) approaches that explicitly incorporate uncertainty into attention mechanisms. This structure clarifies how UAT-Lite differs from and complements existing techniques.
G.1 Calibration and Uncertainty
Despite high predictive accuracy, modern neural networks are often poorly calibrated (Guo et al., 2017). Post-hoc calibration methods such as temperature scaling (Guo et al., 2017), Platt scaling (Platt et al., 1999), and histogram binning (Zadrozny and Elkan, 2001) adjust output confidence without altering internal representations. While simple and effective, these methods are limited by the quality of learned features and do not affect model reasoning.
Ensemble-based approaches improve calibration by averaging predictions from multiple models. Deep Ensembles (Lakshminarayanan et al., 2017) and snapshot ensembles (Huang et al., 2017) are among the strongest calibration baselines, but require training, storing, and evaluating multiple model replicas, resulting in substantial computational and memory overhead. More efficient approximations, such as ensembling pruned attention heads within a single Transformer (Gabetni et al., 2025), reduce cost but still rely on architectural modification or fused ensemble structures. Our work instead targets competitive calibration using inference-time stochasticity within a single pretrained model.
G.2 Bayesian and Stochastic Methods
Bayesian neural networks provide principled uncertainty estimates by placing distributions over model parameters (MacKay, 1992; Neal, 2012). However, scalable approximations such as variational inference (Blundell et al., 2015; Louizos and Welling, 2017) and Laplace methods (Ritter et al., 2018) scale poorly to Transformer-sized models. Bayesian Transformer variants (Xue et al., 2021; Fan et al., 2020) further require architectural modification and retraining, limiting compatibility with pretrained backbones.
Several recent methods estimate uncertainty in weight space. Stochastic Weight Averaging and its Gaussian extension (SWAG) approximate posterior distributions during fine-tuning and improve uncertainty quality in NLI tasks (Talman et al., 2023). Parameter-efficient Bayesian adaptation methods such as C-LoRA (Rahmati et al., 2025) introduce latent variables within low-rank adapters, but require Bayesian fine-tuning and task-specific parameter updates. Other approaches improve uncertainty estimation through post-hoc output heads (Zabolotnyi et al., 2025) or uncertainty-aware loss reweighting during alignment (Wang et al., 2024). These methods primarily influence training dynamics or output distributions and do not modify internal Transformer computations at inference time.
Monte Carlo Dropout (Gal and Ghahramani, 2016) provides a lightweight alternative by approximating Bayesian inference through stochastic forward passes. Standard MC Dropout applies uniform stochasticity across layers and treats uncertainty only as an output-level signal. Our work builds on MC Dropout, but explicitly uses the resulting uncertainty to influence attention during inference.
G.3 Uncertainty-Aware Attention
Calibration and uncertainty issues are particularly pronounced in Transformer-based models. Larger Transformers are often increasingly overconfident (Desai and Durrett, 2020), and semantic uncertainty plays a significant role in language understanding and generation (Kuhn et al., 2023). Most existing methods treat uncertainty as a post-hoc quantity, leaving attention patterns and internal representations unchanged.
Several approaches inject uncertainty directly into attention mechanisms. Bayesian Attention Belief Networks (Zhang et al., 2021) model attention weights as latent random variables, while Sparse Gaussian Process Attention (Chen and Li, 2023) reformulates self-attention using kernel-based Gaussian process inference. Although principled, these methods require architectural changes and variational training, limiting their applicability to pretrained models. Domain-specific uncertainty-aware Transformers, such as TUnA (Ko et al., 2024), introduce uncertainty primarily through output layers rather than through attention modulation itself.
Recent work explores efficient uncertainty mechanisms inspired by ensembles or parameter-efficient adaptation. Hydra Ensembles (Gabetni et al., 2025) approximate Deep Ensembles by fusing multiple attention heads, achieving competitive calibration with reduced cost. However, uncertainty arises from architectural fusion and remains an output-level aggregation. Similarly, C-LoRA (Rahmati et al., 2025) requires Bayesian fine-tuning to model uncertainty. Other methods leverage uncertainty for downstream decision-making, such as selective prediction (Vazhentsev et al., 2025), but do not affect internal attention computation.
Positioning of UAT-Lite.
UAT-Lite integrates epistemic uncertainty into self-attention at inference time, without modifying pretrained weights, training objectives, or model architecture. Unlike post-hoc scalers (e.g., temperature scaling) that adjust confidence only at the output layer, our method uses uncertainty estimates to modulate internal token interactions, aiming to improve reliability under ambiguity, distribution shift, and selective decision-making. Importantly, UAT-Lite is complementary to post-hoc scaling: temperature scaling can be applied on top of our uncertainty-aware inference when the goal is to minimize in-domain calibration error while preserving uncertainty-aware internal computation.
Appendix H AI Usage Statement
We used large language models as auxiliary tools during manuscript preparation. Claude was used for limited coding assistance, including refactoring and debugging support. ChatGPT was used for writing assistance, including grammar checking, spelling correction, and stylistic refinement. All core research contributions, including the scientific ideas, experimental design, implementation, empirical evaluation, analysis, and conclusions, were developed by the authors. All results are based on our own experiments using publicly available datasets and verified code. No AI system was used to generate data, conduct experiments, or draw scientific conclusions.

### v2

@@
 Abstract
-Neural NLP models are often miscalibrated, assigning high confidence to incorrect predictions, which undermines selective prediction and high-stakes deployment.
+Neural NLP models are often miscalibrated and overconfident, assigning high confidence to incorrect predictions and failing to express uncertainty during internal evidence aggregation.
+This undermines selective prediction and high-stakes deployment.
 Post-hoc calibration methods adjust output probabilities but leave internal computation unchanged, while ensemble and Bayesian approaches improve uncertainty at substantial training or storage cost.
-We propose UAT-Lite, an inference-time framework that makes self-attention uncertainty-aware using approximate Bayesian inference via Monte Carlo dropout in pretrained transformer classifiers.
+We propose UAT-Lite, an inference-time framework that makes self-attention uncertainty-aware via Monte Carlo dropout in pretrained transformer classifiers.
+Unlike output-level calibration (e.g., TS), UAT-Lite injects epistemic uncertainty directly into attention, enabling uncertainty-aware routing during contextualization and token-level diagnostic signals beyond global logit rescaling.
 Token-level epistemic uncertainty is estimated from stochastic forward passes and used to modulate self-attention during contextualization, without modifying pretrained weights or training objectives.
 We additionally introduce a layer-wise variance decomposition to diagnose how predictive uncertainty accumulates across transformer depth.
-Across the SQuAD 2.0 answerability, MNLI, and SST-2, UAT-Lite reduces Expected Calibration Error by approximately 20% on average relative to a fine-tuned BERT-base baseline while preserving task accuracy, and improves selective prediction and robustness under distribution shift.
-111github.com/eliashossain001/uq_decomposition
-UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers
+Across SQuAD 2.0 answerability, MNLI, and SST-2, UAT-Lite achieves an average relative ECE reduction of approximately 20% compared with a fine-tuned BERT-base baseline while preserving accuracy, and yields more informative uncertainty behavior for selective prediction under distribution shift.
+111github.com/eliashossain001/uq_decompositionUAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers
 Elias Hossain1††thanks: Corresponding author: mdelias.hossain@ucf.edu, Shubhashis Roy Dipta2, Subash Neupane3, Rajib Rana4, Ravid Shwartz-Ziv5, Ivan Garibay1, Niloofar Yousefi1 1University of Central Florida 2University of Maryland, Baltimore County 3Meharry Medical College 4University of Southern Queensland 5New York University mdelias.hossain@ucf.edu
 1 Introduction
-Pretrained transformer-based language models have achieved state-of-the-art performance across a wide range of NLP tasks, from question answering to clinical decision support Devlin et al.
+Pretrained transformer-based language models achieve state-of-the-art performance across a wide range of NLP tasks, from question answering to clinical decision support Devlin et al.
 (2019); Brown et al.
 (2020).
-Despite these advances, modern transformers exhibit systematic miscalibration, frequently assigning high confidence to incorrect predictions Desai and Durrett (2020).
-This misalignment between confidence and correctness is particularly problematic for selective prediction and related high-stakes applications.
-Fig.˜1 summarizes existing approaches to uncertainty estimation.
-Standard transformer inference is deterministic and does not represent uncertainty during internal computation (Fig.˜1A).
-Post-hoc calibration methods, such as temperature scaling, adjust confidence only after representations are formed, leaving internal reasoning unchanged (Fig.˜1B; Guo et al., 2017).
-Stochastic approaches, including Monte Carlo (MC) dropout and deep ensembles (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017), improve uncertainty estimates but largely treat uncertainty as an output-level signal, without influencing attention.
-This limitation reflects a broader gap: transformers are trained to optimize predictive accuracy rather than confidence reliability Guo et al.
+Despite these advances, modern transformers exhibit systematic miscalibration, often assigning high confidence to incorrect predictions Desai and Durrett (2020).
+This mismatch between confidence and correctness is particularly problematic for selective prediction and other high-stakes settings.
+Fig.˜1 summarizes common approaches to uncertainty estimation.
+Standard transformer inference is deterministic and does not represent uncertainty in internal computation (Fig.˜1A).
+Post-hoc calibration methods, such as temperature scaling (TS), adjust confidence only after representations are formed, leaving internal computation unchanged (Fig.˜1B; Guo et al., 2017).
+Stochastic approaches, including Monte Carlo (MC) dropout and deep ensembles (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017), improve uncertainty estimates but typically treat uncertainty as an output-level signal without influencing attention.
+This gap reflects a broader issue: transformers are trained to optimize predictive accuracy rather than confidence reliability Guo et al.
 (2017).
 While post-hoc methods are lightweight and effective in-domain, they do not alter token interactions.
-Bayesian approaches integrate uncertainty more deeply but typically require architectural modification or specialized training, limiting compatibility with pretrained models.
-This raises the following question: can epistemic uncertainty shape a transformer’s internal attention at inference time, without retraining or modifying pretrained weights?
-We argue that effective uncertainty quantification requires uncertainty to actively shape attention, rather than merely annotate final predictions.
-For ambiguous inputs, a calibrated model should hedge internal reasoning instead of propagating early attention decisions deterministically.
-Existing stochastic inference methods expose uncertainty only at the output level and do not use it to modulate internal attention patterns (Fig.˜1C).
-Motivated by this observation, we propose UAT-Lite, an inference-time framework for encoder-based pretrained transformer classifiers.
-UAT-Lite estimates token-level epistemic uncertainty via Monte Carlo sampling and uses it to downweight attention contributions from unstable tokens, without modifying pretrained weights or training objectives.
-Our main contributions are:
-Uncertainty-Weighted Attention. An inference-time mechanism that modulates self-attention using token-level epistemic uncertainty estimated via Monte Carlo sampling.
-Layer-Wise Uncertainty Attribution. A diagnostic variance decomposition that attributes predictive uncertainty across transformer depth.
-Comprehensive Evaluation.
-Empirical evaluation across multiple NLP benchmarks demonstrating improved calibration, selective prediction, and robustness under distribution shift without degrading task performance.
-UAT-Lite introduces no additional trainable parameters and operates entirely at inference time, requiring only a small number of stochastic forward passes.
-Post-hoc calibration methods such as temperature scaling are complementary rather than competing.
-The framework trades modest inference-time cost for improved reliability while avoiding the training and storage overhead of ensemble-based approaches.
+Bayesian approaches integrate uncertainty more deeply but often require architectural changes or specialized training, which limits compatibility with pretrained models.
+This motivates a central question: can epistemic uncertainty shape a transformer’s attention at inference time without retraining or modifying pretrained weights?
+We argue that effective uncertainty quantification should allow uncertainty to shape attention, not merely annotate final predictions.
+For ambiguous inputs, a reliable model should modulate evidence aggregation rather than propagate early attention decisions deterministically.
+Existing stochastic inference methods expose uncertainty mainly at the output level and do not use it to modulate internal attention patterns (Fig.˜1C).
+Although temperature scaling is highly effective for output calibration, it only rescales logits and cannot provide token-level epistemic signals or uncertainty-aware routing inside the model.
+Motivated by these observations, we propose UAT-Lite, an inference-time framework for encoder-based pretrained transformer classifiers.
+UAT-Lite estimates token-level epistemic uncertainty via MC dropout and uses it to downweight attention contributions from unstable tokens during contextualization, without modifying pretrained weights or training objectives.
+Our main contributions are as follows:
+Uncertainty-Weighted Attention.
+An inference-time mechanism that injects token-level epistemic uncertainty, estimated via MC dropout, into self-attention, downweighting unstable token contributions during contextualization.
+Layer-Wise Uncertainty Attribution.
+A variance decomposition that attributes predictive uncertainty across transformer depth, enabling diagnostic analysis of where uncertainty amplifies during inference.
+Comprehensive Evaluation. Experiments on SQuAD 2.0 answerability, MNLI, and SST-2 that show improved calibration and selective prediction, along with robustness analyses under distribution shift.
+UAT-Lite introduces no additional trainable parameters and operates strictly at inference time, using a controllable Monte Carlo budget (e.g., ) to trade compute for reliability.
+UAT-Lite is not intended to replace lightweight output-level calibrators such as TS, which is optimized on development logits and often performs best in-domain.
+While TS rescales final logits after representations are formed, UAT-Lite modulates internal evidence aggregation by making attention uncertainty-aware.
+Concretely, UAT-Lite provides (i) token-level epistemic signals, (ii) uncertainty-aware routing via attention modulation, and (iii) layer-wise diagnostic attribution; TS provides only output probability rescaling.
+Because the two mechanisms act at different stages, they are complementary and can be stacked (UAT-Lite +TS) when calibrated probabilities are needed.
+Overall, the framework trades a controllable inference-time cost for improved reliability while avoiding the training and storage overhead of ensemble-based alternatives.
 Our study focuses on encoder-based transformers; extending uncertainty-aware attention to decoder-only generative models is left for future work.
+2 Related Work
+Uncertainty estimation and calibration for neural networks have been studied extensively.
+Post-hoc methods such as temperature scaling and Platt scaling (Guo et al., 2017; Platt et al., 1999) provide simple and effective confidence correction without altering model internals.
+Ensemble-based approaches, including Deep Ensembles (Lakshminarayanan et al., 2017), often yield strong calibration performance but incur substantial training and storage overhead.
+Bayesian neural networks, stochastic inference methods, and recent Bayesian transformer variants (Blundell et al., 2015; Gal and Ghahramani, 2016; Ritter et al., 2018; Fan et al., 2020; Zhang et al., 2021; Chen and Li, 2023) offer principled approaches to uncertainty modeling, but they typically require architectural modifications, variational training, or task-specific retraining.
+By contrast, UAT-Lite integrates epistemic uncertainty directly into self-attention at inference time, preserving pretrained weights and standard training pipelines while enabling uncertainty-aware internal computation.
+A more comprehensive discussion of related work and detailed comparisons is provided in App.˜G.
 3 Preliminaries
-Bayesian inference and uncertainty.
-Bayesian analysis incorporates parameter uncertainty by averaging predictions over probable parameter values inferred from the training data.
-This perspective distinguishes epistemic uncertainty, which arises from limited data or insufficient coverage and reflects uncertainty about the model itself, from aleatoric uncertainty, which represents irreducible noise in the observations.
-In large neural models, epistemic uncertainty becomes particularly significant in the presence of semantic ambiguity or distributional shifts, where confident yet incorrect predictions frequently occur.
-Monte Carlo (MC) Dropout.
-We briefly recall MC Dropout as an approximate Bayesian inference technique used throughout this work.
-MC dropout retains dropout layers at inference time and performs multiple stochastic forward passes, yielding a predictive distribution whose variance estimates epistemic uncertainty and whose mean defines the final prediction.
-Conventional MC dropout treats uncertainty solely at the output level and does not influence internal representations or attention mechanisms, motivating methods that integrate uncertainty directly into the inference process.
-Token-level uncertainty as a proxy signal.
-While epistemic uncertainty is often measured at the output level, internal transformer computation is structured around token-level representations and attention interactions.
-Consequently, uncertainty estimates derived from stochastic embedding representations can serve as a lightweight proxy for identifying unstable or ambiguous token contributions.
-This perspective motivates the use of token-level uncertainty as a control signal for modulating attention during inference, without requiring architectural changes or retraining.
-3 Related Work
-Uncertainty estimation and calibration for neural networks have been widely studied, with post-hoc methods such as temperature scaling and Platt scaling (Guo et al., 2017; Platt et al., 1999) providing simple and effective confidence correction without altering model internals.
-Ensemble-based approaches, including Deep Ensembles (Lakshminarayanan et al., 2017), yield strong calibration but incur substantial training and storage overhead.
-Bayesian neural networks, stochastic inference methods, and recent Bayesian Transformer variants (Blundell et al., 2015; Gal and Ghahramani, 2016; Ritter et al., 2018; Fan et al., 2020; Zhang et al., 2021; Chen and Li, 2023) offer principled uncertainty modeling but typically require architectural modification, variational training, or task-specific retraining.
-In contrast, UAT-Lite integrates epistemic uncertainty directly into self-attention at inference time, preserving pretrained weights and standard training pipelines while enabling uncertainty-aware internal computation.
-A comprehensive discussion of related work and detailed comparisons is provided in App.˜G.
+We briefly review the uncertainty concepts and stochastic inference protocol used throughout the paper.
+Bayesian perspectives distinguish epistemic uncertainty, which arises from limited data or distribution shift, from aleatoric uncertainty, which reflects irreducible observation noise.
+In our setting, epistemic effects dominate under semantic ambiguity and distribution shift, where transformers can be confidently wrong.
+We use MC dropout as a lightweight approximate Bayesian inference method.
+Specifically, dropout is retained at test time, and we run stochastic forward passes to obtain a predictive distribution whose mean is used for prediction and whose variability serves as an estimate of epistemic uncertainty.
+Standard MC dropout treats uncertainty as an output-level signal and does not influence internal computation.
+Because transformer inference is organized around token representations and attention interactions, we use token-level uncertainty derived from stochastic embeddings as a proxy for unstable token contributions.
+This motivates modulating attention with token uncertainty at inference time, without retraining or changing pretrained weights.
 4 UAT-Lite
-UAT-Lite is an uncertainty-aware extension of pretrained transformer classifiers designed to improve confidence calibration at inference time without modifying model architecture, training objectives, or learned parameters.
-The framework integrates approximate Bayesian uncertainty estimation into internal attention computation, allowing epistemic uncertainty to actively influence contextualization while preserving the base transformer’s parameterization.
-In addition, UAT-Lite introduces a diagnostic layer-wise variance decomposition that characterizes how predictive uncertainty evolves across transformer depth, enabling post-hoc analysis without affecting the forward pass.
-Final predictions are produced via stochastic inference with confidence-aware decision shaping.
-Overall, UAT-Lite preserves the training cost and parameter count of the underlying encoder, incurring only modest inference-time overhead.
+UAT-Lite is an inference-time extension of pretrained transformer classifiers that injects epistemic uncertainty into self-attention to improve confidence reliability without modifying the model architecture, training objectives, or learned parameters.
+The framework uses MC dropout to estimate token-level variability and to modulate attention during contextualization.
+UAT-Lite also provides a diagnostic layer-wise variance decomposition that characterizes how predictive uncertainty accumulates across transformer depth without affecting the forward pass.
+Final predictions are produced by aggregating stochastic logits, and optional TS can be applied to the MC-mean logits when calibrated probabilities are required.
 4.1 Problem Setup and Notation
 We consider an encoder-based pretrained transformer for classification.
-Given an input sequence , the model maps through stacked layers to contextual representations , which are converted to a predictive distribution by a task-specific classification head.
-In standard evaluation mode, pretrained transformers produce a single point estimate: dropout is disabled and the same input yields the same representations and prediction given fixed parameters.
-While effective for accuracy, this design provides no mechanism for epistemic uncertainty.
-Our goal is to introduce controlled inference-time stochasticity that allows uncertainty to influence internal attention without modifying training or parameters.
+Given an input sequence , the model maps through layers to contextual representations , which are converted to a predictive distribution by a task-specific head.
+In standard evaluation, inference is deterministic: dropout is disabled, and the same input yields the same prediction for fixed parameters.
+While effective for accuracy, this setting provides no mechanism for representing epistemic uncertainty during internal computation.
+Our goal is to introduce controlled inference-time stochasticity and use it to modulate attention without retraining or parameter updates.
 4.2 Estimating Uncertainty via MC Sampling
 We estimate predictive uncertainty using MC dropout, a lightweight approximation to Bayesian inference.
-Dropout layers are retained during inference and stochastic forward passes are performed, inducing variability via random neuron masking.
-Aggregating these samples yields a predictive mean and an estimate of epistemic variability across stochastic passes during evaluation.
+Dropout is retained at test time, and we perform stochastic forward passes, each induced by a different dropout mask.
+Aggregating the resulting logits yields a predictive mean and an estimate of epistemic variability across passes.
 Uncertainty type.
 Following Kendall and Gal (2017), predictive uncertainty decomposes into aleatoric and epistemic components.
-MC dropout primarily captures epistemic uncertainty through approximate posterior sampling during inference.
+MC dropout primarily captures epistemic uncertainty through approximate posterior sampling at inference time.
 Component-specific dropout.
-To balance representational stability and uncertainty estimation, we apply component-specific dropout during inference (0.1 for embeddings, 0.2 for attention layers, and 0.3 for feed-forward layers).
-These rates follow common practice in transformer fine-tuning and are kept fixed across tasks; calibration trends are stable under moderate variations of these values.
+To balance variability and stability, we use component-specific dropout rates during inference: 0.1 for embeddings, 0.2 for attention layers, and 0.3 for feed-forward layers.
+These rates are fixed across tasks, and calibration trends remain stable under moderate rate variations.
 Single-pass sampling protocol.
-We use a single-pass MC inference procedure in which the same stochastic forward passes are used both to (i) estimate token-level embedding uncertainty and (ii) obtain the final MC predictive distribution.
-As detailed in Algorithm 1, for each pass we sample dropout masks, compute embeddings , propagate the input through the encoder with uncertainty-weighted attention, and record logits for aggregation.
-Token-level uncertainty is computed by aggregating embedding variability across these passes and is reused consistently within the same inference run.
-This design avoids any separate embedding-only sampling stage and ensures that inference-time overhead scales linearly with .
+We use a single-pass MC procedure in which the same stochastic forward passes are used to (i) estimate token-level embedding uncertainty and (ii) compute the final MC predictive distribution.
+Concretely, during pass we sample dropout masks and compute embeddings , then update running uncertainty estimates from the embedding samples observed so far.
+To avoid circular dependence within a pass, we modulate attention in pass using the lagged estimate , with , as summarized in Algorithm 1.
+After passes, the final token uncertainty is , and predictions are obtained by aggregating logits .
+This design avoids a separate embedding-only stage and ensures that overhead scales linearly in .
 4.3 Uncertainty-Weighted Self-Attention
-We incorporate epistemic uncertainty directly into self-attention by modulating attention logits using token-level uncertainty estimates computed at the embedding layer during contextualization.
+We incorporate epistemic uncertainty into self-attention by modulating attention logits using token-level uncertainty estimated at the embedding layer.
 Token-level uncertainty estimation.
-For each input token , we obtain stochastic embedding samples . Token-level uncertainty is defined as
-where denotes the embedding dimensionality. We interpret as an embedding-level proxy for epistemic uncertainty induced by dropout.
-This quantity is computed once per input and shared across all layers and attention heads.
-The computational cost is negligible relative to Monte Carlo methods, as it involves only per-token operations over the sampled embeddings and does not require additional forward passes.
-Scale sensitivity.
-The magnitude of remains well-behaved due to embedding normalization and averaging across dimensions.
-Sensitivity analysis over and (§˜D.1) shows stable calibration across a broad range of values; explicit normalization is therefore unnecessary.
+For each token , we obtain stochastic embedding samples via MC dropout. We define token uncertainty as
+where is the embedding dimension. We treat as an embedding-level proxy for the epistemic uncertainty induced by dropout and use it as a control signal for uncertainty-aware attention routing.
+Single-pass / online estimation.
+To avoid a separate embedding-only sampling stage, we estimate online during the same stochastic passes used for prediction.
+Specifically, after completing pass , we maintain a running estimate (e.g., via Welford updates), and we use to modulate attention in pass .
+After passes, the final estimate aggregates variability across all samples.
+During the MC run, each pass uses the lagged estimate consistently across all layers and heads.
+Scale/geometry diagnostics.
+We test whether is a trivial artifact of embedding scale or geometry through embedding-rescaling invariance, correlations with output-space epistemic measures (entropy/MI), and uncertainty-definition ablations (§˜D.8).
 Attention modulation.
-Let , , and denote query, key, and value projections. Standard scaled dot-product attention logits are
-We modulate these logits using token-level uncertainty:
-where is a fixed uncertainty penalty. Attention probabilities are obtained via softmax:
+Let , , and denote the query, key, and value projections. Standard scaled dot-product attention logits are
+We define uncertainty-aware logits by attenuating the pre-softmax scores:
+where is a fixed penalty parameter and specifies the projection variant.
+Positioning vs.
+TS.
+UAT-Lite does not replace TS: TS rescales final logits post hoc, whereas UAT-Lite modulates pre-softmax attention logits using token-level epistemic uncertainty, enabling uncertainty-aware routing and token-level diagnostics beyond output rescaling.
+We analyze sensitivity to and compare penalty forms (exponential vs.
+linear vs.
+reciprocal) in §˜D.6.
+Projection variants (Q/K/V).
+Using , we instantiate as follows:
+We then obtain attention probabilities via the softmax:
+We additionally consider V-only and QKV by applying the same factor to the value path, , and using as the attention output.
 Relation to prior work.
-Design choices.
-Uncertainty modulation is applied uniformly across layers and heads and affects only the key dimension in the attention mechanism.
-Unless otherwise stated, the uncertainty penalty (default ) is set once on development data and held fixed across tasks.
-§˜D.1 shows that calibration is stable across a broad range of and .
-The method introduces no additional trainable parameters beyond the base model.
+Prior uncertainty-aware attention methods typically require architectural modification or variational training (Chen and Li, 2023; Zhang et al., 2021), whereas our approach operates entirely at inference time.
+Unless otherwise stated, we default to Q-only, which empirically provides a favorable calibration–accuracy trade-off in our experiments in Section 6.5.
 4.4 Layer-Wise Uncertainty Attribution
-We analyze how predictive uncertainty propagates across transformer depth under inference-time stochasticity. This analysis is diagnostic and does not affect forward computation or model behavior.
+We analyze how predictive uncertainty propagates across transformer depth under inference-time stochasticity. This analysis is purely diagnostic and does not alter model behavior.
 Layer-wise variance decomposition.
-Using the law of total variance, total predictive variance can be decomposed into layer-specific contributions under inference-time stochasticity. A formal derivation is provided in App.˜A.
-Theorem 1 (Layer-Wise Variance Decomposition).
-Under inference-time stochasticity, total predictive variance can be expressed as the sum of layer-specific variance contributions.
+Using the law of total variance, we decompose the predictive variance under MC dropout into layer-indexed contributions (App.˜A).
+Proposition 1 (Diagnostic Layer-Wise Variance Decomposition).
+Under the stochastic inference procedure used in this work, predictive variance admits a recursive decomposition across transformer depth via repeated application of the law of total variance.
 Scope and interpretation.
-This result follows from standard variance identities and is not a novel mathematical contribution.
-Its value lies in enabling diagnostic inspection of uncertainty evolution across transformer depth, which is unavailable in post-hoc calibration methods.
-Layer-wise contributions should be interpreted as approximate signals rather than strictly causal attributions.
+This decomposition follows from standard variance identities and is not claimed to be a novel theoretical result.
+Its purpose is to diagnose where uncertainty concentrates or amplifies across depth.
+The resulting contributions should therefore be interpreted as approximate diagnostic signals rather than causal attributions.
 4.5 Confidence-Aware Decision Shaping
-MC dropout produces a distribution of stochastic predictions.
-The predictive mean is computed from the stochastic forward passes.
-Post-hoc temperature scaling is applied to the Monte Carlo mean logits, with the temperature parameter learned on the validation set to calibrate predictive confidence.
+MC dropout yields logits from stochastic passes.
+We compute the MC-mean logits as and derive the predictive distribution from .
+When used, post-hoc temperature scaling is fit on the validation set and applied to at test time.
 Selective prediction.
-Uncertainty estimates enable abstention via confidence thresholds selected on the validation split of each dataset and applied uniformly at test time across all evaluated methods.
+We perform abstention by thresholding confidence on the validation split and applying the same thresholds at test time for all methods.
 5 Experimental Setup
 5.1 Datasets
-We evaluate on five benchmarks spanning general NLP and clinical domains commonly used in prior calibration and robustness studies.
+We evaluate on five benchmarks spanning general-domain NLP and clinical-language settings that are commonly used in prior calibration and robustness studies.
 All experiments focus exclusively on encoder-based pretrained transformer classifiers (BERT-family backbones), rather than decoder-only language models.
-General NLP datasets include SQuAD 2.0 Rajpurkar et al.
-(2018) (QA with unanswerable questions), MNLI Williams et al.
-(2018) (natural language inference with matched/mismatched splits for distribution shift analysis), and SST-2 Socher et al.
-(2013) (binary sentiment classification).
-Clinical benchmarks include MedQA Jin et al.
+The general-domain datasets include SQuAD 2.0 Rajpurkar et al.
+(2018), a question-answering benchmark with unanswerable questions; MNLI Williams et al.
+(2018), a natural language inference benchmark with matched and mismatched splits that support distribution-shift analysis; and SST-2 Socher et al.
+(2013), a binary sentiment classification benchmark.
+The clinical-domain benchmarks include MedQA Jin et al.
 (2021) and PubMedQA Jin et al.
-(2019), which are used to assess domain transfer rather than clinical validity.
+(2019), which we use to assess domain transfer rather than clinical validity.
 5.2 Baselines, Metrics, and Training Overview
-We evaluate our approach against BERT-base, MC Dropout, Temperature Scaling (TS), Deep Ensembles, and representative Bayesian last-layer methods.
-All models share the same pretrained backbone, fine-tuning recipe, and evaluation protocol; detailed baseline configurations are described in App.˜C.
-Calibration is evaluated using Expected Calibration Error (ECE), alongside robustness under distribution shift and selective prediction metrics.
-Results are reported as mean standard deviation over five random seeds, with deterministic baselines exhibiting negligible variance.
+We compare our approach against BERT-base, MC Dropout, TS, Deep Ensembles, and representative Bayesian last-layer methods.
+All models share the same pretrained backbone, fine-tuning recipe, and evaluation protocol; detailed baseline configurations are provided in App.˜C.
+Calibration is evaluated using Expected Calibration Error (ECE), together with robustness under distribution shift and selective prediction metrics.
+Results are reported as mean standard deviation over five random seeds, while deterministic baselines exhibit negligible variance.
 The proposed method introduces no additional trainable parameters and operates entirely at inference time.
+Negation diagnostic.
+We include a controlled negation perturbation as a diagnostic of uncertainty responsiveness under semantic destabilization.
+Because negation may alter the true label, we emphasize label-free stability signals (e.g., Margin and Entropy) rather than treating this setting as a robustness test of correctness.
+Full details and examples are provided in §˜D.9.
 Uncertainty penalty.
-Unless otherwise stated, we use a single uncertainty penalty , set once on development data and held fixed across all tasks and datasets.
-§˜D.1 shows that calibration performance is stable across a broad range of and Monte Carlo budgets , indicating that the method does not rely on sensitive hyperparameter tuning in practice across the evaluated settings.
+Unless otherwise stated, we use a single uncertainty penalty , which is selected once on development data and then held fixed across all tasks and datasets.
+§˜D.11 shows that calibration performance remains stable across a broad range of values and Monte Carlo budgets , indicating that the method does not rely on sensitive hyperparameter tuning in practice for the evaluated settings.
+In particular, across and , ECE varies within a narrow band (range ), while accuracy changes are negligible.
+Increasing improves calibration, with diminishing returns beyond , suggesting that stable behavior does not require expensive Monte Carlo budgets or fine-grained tuning.
 6 Results
 6.1 General NLP Calibration Performance
-Table 1 reports multi-seed calibration results (mean standard deviation over five seeds), computed using identical ECE binning and confidence definitions across methods.
-Deterministic baselines exhibit negligible variance, while stochastic methods show small but consistent variability.
-UAT-Lite improves calibration relative to the unscaled fine-tuned BERT baseline, reducing average ECE from 0.117 to 0.094 (19.9% relative reduction).
-The largest gain is observed on MNLI (0.110 0.067), a task characterized by semantic ambiguity, where uncertainty-aware attention is particularly beneficial.
-These improvements persist relative to both uniform and component-specific MC Dropout, indicating that the gains arise from uncertainty-weighted attention rather than from stochasticity introduced by sampling.
-Temperature scaling (TS) achieves the lowest in-domain ECE across all tasks, as expected for a post-hoc method explicitly optimized to minimize calibration error on held-out validation data.
-Our objective is complementary rather than competitive: while TS rescales output probabilities, it does not alter the internal representations or attention patterns that produce a prediction.
-In contrast, UAT-Lite introduces uncertainty-aware internal computation at inference time, allowing epistemic uncertainty to influence token interactions during contextualization.
-Because TS operates purely at the output level, it can be directly stacked on top of UAT-Lite without modification; we focus here on isolating the effects of uncertainty-aware attention and evaluate its benefits under distribution shift and selective prediction in subsequent sections.
-Among Bayesian-style alternatives, SNGP achieves lower ECE but requires architectural modification and specialized training.
-Deep Ensembles provide a strong reference point (Avg ECE 0.089), but incur substantial training and storage overhead due to maintaining multiple model replicas.
-In contrast, UAT-Lite operates on a single pretrained transformer and introduces uncertainty handling entirely at inference time.
-As a result, it substantially narrows the calibration gap to Deep Ensembles while preserving deployment efficiency.
-| Method | SQuAD (ECE ) | MNLI (ECE ) | SST-2 (ECE ) | Avg (ECE ) |
-| BERT-base | 0.186 0.000 | 0.110 0.000 | 0.055 0.000 | 0.117 |
-| MC Dropout (uniform ) | 0.173 0.001 | 0.090 0.001 | 0.039 0.001 | 0.101 |
-| MC Dropout (component-specific) | 0.186 0.000 | 0.109 0.001 | 0.054 0.001 | 0.116 |
-| Temp. Scaling | 0.057 0.000 | 0.028 0.000 | 0.029 0.000 | 0.038 |
-| SNGP | 0.001 0.000 | 0.021 0.000 | 0.009 0.000 | 0.010 |
-| VI-LastLayer | 0.119 0.001 | 0.096 0.000 | 0.136 0.002 | 0.117 |
-| Deep Ensemble (5) | 0.168 0.000 | 0.060 0.000 | 0.037 0.000 | 0.089 |
-| UAT-Lite | 0.163 0.002 | 0.067 0.001 | 0.051 0.004 | 0.094 |
-| vs. Baseline | 12.5% | 39.0% | 6.6% | 19.9% |
+Table 1 reports multi-seed calibration results on SQuAD, MNLI, and SST-2 (ECE ; mean standard deviation over five seeds) under a shared evaluation protocol.
+Across tasks, UAT-Lite improves calibration relative to the unscaled fine-tuned BERT baseline (e.g., average ECE decreases from to ), with the largest improvement on MNLI ().
+It also outperforms MC Dropout baselines, including uniform and component-specific variants, suggesting that the gains are not explained by stochasticity alone.
+Fig.
+4 shows that global- or document-level MC Dropout improves over the deterministic baseline but remains behind TS, which achieves the lowest in-domain ECE.
+We therefore view UAT-Lite as complementary to TS: UAT-Lite modulates internal evidence aggregation, whereas TS rescales output logits.
+The bottom panel of Table 1 reports MedQA and PubMedQA (single seed; accuracy reported in separate columns) as a clinical domain-transfer stress test.
+MedQA analysis by linguistic bucket (negation, numeric expressions, biomedical terminology, and hedging markers; Table 12) shows a consistent pattern: UAT-Lite improves accuracy across all phenomena, while UAT-Lite +TS attains the lowest calibration error in every bucket.
+Overall, TS is strongest for in-domain probability calibration, UAT-Lite provides complementary uncertainty-aware inference, and UAT-Lite +TS yields the best trade-off in our evaluations.
+Panel A: General NLP (multi-seed) Method SQuAD (ECE ) MNLI (ECE ) SST-2 (ECE ) Avg (ECE ) BERT-base 0.1868 0.000 0.0816 0.000 0.0531 0.000 0.1072 Base+TS 0.0788 0.000 0.0148 0.000 0.0163 0.000 0.0366 MC Dropout (uniform ) 0.173 0.001 0.090 0.001 0.039 0.001 0.101 MC Dropout (component-specific) 0.186 0.000 0.109 0.001 0.054 0.001 0.116 SNGP 0.001 0.000 0.021 0.000 0.009 0.000 0.010 VI-LastLayer 0.119 0.001 0.096 0.000 0.136 0.002 0.117 Deep Ensemble (5) 0.168 0.000 0.060 0.000 0.037 0.000 0.089 Global MC Dropout (M=10) 0.1600 0.000 0.0565 0.000 0.0501 0.000 0.0889 UAT-Lite (M=10) 0.1647 0.000 0.0638 0.000 0.0608 0.000 0.0964 UAT-Lite+TS (M=10) 0.0760 0.000 0.0201 0.000 0.0191 0.000 0.0384 Panel B: Clinical QA (single seed) Method MedQA (ECE ) MedQA (Acc ) PubMedQA (ECE ) PubMedQA (Acc ) Baseline 0.033 0.246 0.114 0.620 Base+TS 0.023 0.246 0.128 0.620 Deep Ensemble (5) 0.007 0.268 0.051 0.520 UAT-Lite 0.040 0.249 0.103 0.640 UAT-Lite+TS 0.0069 0.251 0.1018 0.630
 6.2 Distribution Shift Robustness
-We evaluate calibration robustness under distribution shift on MNLI by transferring models fine-tuned on the matched (in-domain) split to the mismatched (out-of-distribution) split.
-Results in Table 2 are reported as mean standard deviation over five seeds, with ECE computed consistently.
-Absolute ECE values differ from the in-domain setting because MNLI shift evaluation is framed as a binary decision (entailment vs.
-non-entailment), and confidence is computed from the Monte Carlo–averaged predictive distribution using the maximum softmax probability.
-This affects the scale of ECE but preserves meaningful relative comparisons across methods.
-BERT-base exhibits clear calibration drift under shift, with ECE increasing from the matched to the mismatched split ().
-In contrast, UAT-Lite yields lower ECE on both in-domain and out-of-domain splits and reduces calibration drift (), resulting in a substantially lower average ECE across the two conditions.
-Importantly, this evaluation isolates the effect of uncertainty-aware internal computation under shift.
-Post-hoc calibration methods such as temperature scaling are intentionally omitted here, as they rescale output confidences without modifying the internal representations that give rise to distributional sensitivity.
-The improved stability observed for UAT-Lite therefore reflects the benefit of allowing epistemic uncertainty to modulate attention during contextualization, rather than relying solely on output-level correction.
-Overall, these results indicate that uncertainty-weighted attention produces more stable confidence estimates under matched-to-mismatched distribution shift, a setting where overconfident errors are particularly costly.
-| Method | ID ECE | OOD ECE | ECE | Avg ECE |
-| BERT-base | 0.2893 0.0000 | 0.3017 0.0000 | 0.0124 | 0.2955 |
-| UAT-Lite | 0.0219 0.0007 | 0.0145 0.0007 | -0.0074 | 0.0182 |
+We evaluate uncertainty under covariate shift in two settings: (i) MNLI transfer (matchedmismatched; mild shift) and (ii) four OOD suites with stronger shifts (HANS, ANLI, SNLI, IMDb).
+Table 2 reports MNLI transfer results (mean standard deviation over five seeds), and Table 5 reports selective prediction on OOD suites, including admitted counts .
+Calibration baselines and composability with TS.
+Temperature scaling (TS) is the strongest post-hoc baseline for reducing marginal ECE via global logit rescaling.
+By contrast, UAT-Lite modulates attention-level evidence aggregation using MC-dropout uncertainty and composes naturally with TS.
+On MNLI transfer, UAT-Lite +TS matches TS-level average ECE with essentially unchanged accuracy (Table 2).
+Selective prediction at fixed confidence thresholds.
+We report Cov@, Acc@, , and AURC for (Table 5).
+Across OOD suites, TS reliably reduces ECE but does not uniformly dominate selective prediction, especially when score compression reduces high- coverage.
+Empirically, UAT-Lite +TS is the most reliable configuration, preserving TS-level calibration while retaining attention-level gating.
+Shortcut-stress test on HANS.
+HANS probes heuristic shortcut reliance.
+We additionally report a per-heuristic breakdown and a non-entailment-only subset (Table 6), which isolates the dominant failure mode of overconfident non-entailment errors.
+On this diagnostic, UAT-Lite and UAT-Lite +TS improve accuracy over Base and Base+TS, suggesting that uncertainty-weighted attention mitigates shortcut-driven evidence beyond output-only rescaling.
+| Method | ID ECE | OOD ECE | ECE | Avg ECE | ID Acc | OOD Acc |
+| Base | 0.08090.0040 | 0.05550.0000 | 0.0040 | 0.06820.0020 | 0.71620.0048 | 0.74590.0000 |
+| Base+TS | 0.01470.0031 | 0.02080.0022 | 0.00610.0040 | 0.01770.0018 | 0.71620.0048 | 0.74590.0000 |
+| Global MC Dropout (M=10) | 0.05860.0037 | 0.03340.0023 | 0.0033 | 0.04600.0026 | 0.70900.0047 | 0.73980.0012 |
+| UAT-LITE (M=10) | 0.06500.0028 | 0.04020.0012 | 0.0038 | 0.05260.0010 | 0.70990.0038 | 0.73950.0014 |
+| UAT-LITE+TS (M=10) | 0.01630.0022 | 0.01990.0013 | 0.00350.0017 | 0.01810.0016 | 0.71050.0047 | 0.73940.0004 |
 6.3 Selective Prediction Performance
-Table 3 reports selective prediction performance using AURC (lower is better), averaged over five seeds.
-All selection criteria are determined on validation data and held fixed across methods.
-Because the Baseline is deterministic, its variance across seeds is negligible, whereas UAT-Lite exhibits small but consistent variability across runs.
-Selective prediction directly evaluates the quality of uncertainty estimates rather than calibration alone: a reliable model should abstain on difficult inputs while maintaining high coverage on confident predictions.
-Accordingly, this evaluation isolates whether uncertainty estimates meaningfully align with prediction risk under identical settings.
-On MNLI, UAT-Lite achieves substantially higher selective coverage, increasing from 0.714 to 0.869, while reducing AURC from 0.164 to 0.053.
-This reflects a markedly improved risk–coverage trade-off, where abstention is applied preferentially to higher-risk inputs rather than uniformly.
-On SQuAD 2.0, coverage increases from 0.501 to 0.690 and AURC decreases from 0.447 to 0.289, indicating more effective abstention behavior on ambiguous or unanswerable inputs.
-These gains demonstrate that uncertainty-weighted attention yields confidence estimates that better reflect underlying uncertainty.
-Post-hoc calibration methods such as temperature scaling are intentionally omitted from this analysis, as they modify output probabilities without altering the confidence ranking that governs abstention behavior.
-These improvements reflect uncertainty-aware internal computation.
-| Dataset | Method | Coverage@0.9 | Coverage@0.8 | Coverage@0.7 | AURC |
-| MNLI | Baseline | 0.7139 0.0000 | 0.7527 0.0000 | 0.7926 0.0000 | 0.1640 0.0000 |
-| MNLI | UAT-Lite | 0.8692 0.0003 | 0.9030 0.0006 | 0.9275 0.0010 | 0.0529 0.0002 |
-| SQuAD 2.0 | Baseline | 0.5007 0.0000 | 0.5071 0.0000 | 0.5050 0.0000 | 0.4472 0.0000 |
-| SQuAD 2.0 | UAT-Lite | 0.6904 0.0006 | 0.7047 0.0009 | 0.7163 0.0010 | 0.2892 0.0007 |
-6.4 Computational Efficiency
-Table 4 reports inference cost and calibration performance on SQuAD 2.0 under a single-accelerator, sequential inference setting.
-We focus on methods with clearly defined deployment behavior and directly measurable inference overhead.
-Predictive accuracy is not reported in this analysis, as all compared methods exhibit comparable task performance on SQuAD 2.0 and the purpose of this section is to isolate calibration behavior under differing inference-time costs.
-UAT-Lite with incurs higher latency than the deterministic baseline (86.31 ms vs.
-17.69 ms), reflecting the cost of repeated stochastic forward passes, and yields a throughput of 13.9 examples/sec.
-Under the same sequential evaluation protocol, this latency is comparable to that of a 5-model Deep Ensemble (88.80 ms), while requiring only a single pretrained model (109M parameters) rather than storing and maintaining multiple independently trained replicas (547M parameters).
-Temperature scaling adds negligible runtime overhead relative to the baseline, but operates purely as an output-level post-hoc calibration method and does not alter internal computation.
-Overall, UAT-Lite incurs additional inference-time cost in exchange for improved calibration, while avoiding the substantial training and storage overhead associated with ensemble-based approaches and preserving a single-model deployment footprint effectively.
-| Method | Params | Samples | Latency (ms) | Memory (GB) | Throughput | ECE |
-| BERT-base | 109M | 1 | 17.69 | 1.78 | 68.9 | 0.186 |
-| Temperature Scaling | 109M | 1 | 17.71 | 1.78 | 69.0 | 0.057 |
-| Deep Ensemble (5) | 547M | 5 | 88.80 | — | — | 0.168 |
-| UAT-Lite | 109M | 10 | 86.31 | 1.81 | 13.9 | 0.163 |
-6.5 Negation-Aware Attention
-Fig.˜3 qualitatively illustrates how uncertainty-weighted attention modulates token interactions under linguistic negation.
-We compare attention patterns from a standard transformer and UAT-Lite under negated linguistic constructions.
-While absolute attention maps remain similar due to softmax normalization, the difference heatmap reveals small but structured redistributions.
-Uncertainty-weighted attention slightly downweights sentiment-bearing tokens (e.g., “positively”) while reallocating attention toward negation cues and predicate structure associated with the phrase “did not respond” in context.
-These changes are intentionally modest and should be interpreted as directional rather than large-scale attention reconfiguration.
-Accordingly, this example is illustrative rather than a precise token-level attribution.
-Corpus-level analysis shows that similar directional trends recur across negated and uncertainty-sensitive sentences (see §˜D.4 for additional qualitative evidence).
+Table 4 reports threshold-based selective prediction results on MNLI and SQuAD 2.0, including TS and global MC Dropout baselines.
+Confidence is defined as .
+For , we report coverage (Cov@), admitted accuracy (Acc@), admitted count (), and AURC (risk–coverage; ).
+Thresholds are selected on validation data for each method and then applied at test time; we also report to make coverage shifts explicit.
+Selective prediction evaluates whether uncertainty aligns with prediction risk: a reliable model should abstain on higher-risk inputs while retaining coverage on lower-risk, high-confidence examples.
+This setting is particularly informative under fixed confidence thresholds, where improved marginal calibration does not necessarily imply better abstention behavior.
+Effect of TS under fixed thresholds.
+We include TS because, although it preserves the predicted label, probability rescaling changes which examples exceed a fixed threshold .
+As a result, TS can substantially alter threshold-level behavior even when accuracy remains unchanged.
+On SQuAD 2.0, for example, TS can markedly reduce Cov@0.9, reflecting confidence compression rather than a reporting artifact.
+Effect of uncertainty-aware inference.
+By contrast, UAT-Lite modifies internal evidence aggregation through uncertainty-aware attention.
+This changes how confidence is formed before thresholding and can improve alignment between confidence and risk at fixed operating points.
+Overall, UAT-Lite, and especially UAT-Lite +TS, yield more reliable risk–coverage behavior, whereas TS alone can substantially shift coverage through score compression; see Table 4 and Table 5.
+6.4 Computational Efficiency and Latency
+Because UAT-Lite relies on Monte Carlo sampling, it incurs additional inference-time cost that scales approximately linearly with the MC budget under a sequential sampling loop (i.e., stochastic forward passes).
+To quantify end-to-end overhead in a single canonical setting, we report wall-clock latency on an NVIDIA A100 GPU with batch size and sequence length .
+UAT-Lite uses stochastic passes implemented as a sequential Monte Carlo loop; under this setting, the deterministic model requires ms per batch, whereas UAT-Lite requires ms, corresponding to a overhead.
+Temperature scaling adds negligible test-time cost ( ms) because it only applies a scalar rescaling of the logits.
+Table 18 reports the resulting overhead.
+At , UAT-Lite incurs a substantial slowdown relative to deterministic inference, whereas TS remains essentially free.
+Accordingly, UAT-Lite is best suited to offline or batch evaluation, or to selectively triggered regimes (e.g., invoked only for low-confidence or high-risk inputs), rather than as an always-on mechanism in latency-critical deployments.
 6.5 Ablation Study
-Fig.˜4 decomposes UAT-Lite into its constituent components to isolate their individual contributions.
-Uncertainty-weighted attention alone reduces ECE from 0.169 to 0.134 (21.1% relative reduction), indicating that attention modulation based on epistemic uncertainty is the primary driver of calibration improvement under controlled comparisons.
-Embedding-level uncertainty alone slightly degrades calibration, suggesting that stochasticity without uncertainty-aware computation is insufficient.
-In contrast, combining uncertainty-weighted attention with confidence-aware decision shaping yields the largest improvement (28.2% relative reduction), exceeding either component in isolation and demonstrating their complementarity.
+Fig.˜5 decomposes UAT-Lite into its constituent components to isolate their individual contributions.
+Uncertainty-weighted attention alone reduces ECE under controlled comparisons, indicating that attention modulation based on epistemic uncertainty is a major driver of the observed calibration gains.
+By contrast, embedding-level stochasticity alone can slightly degrade calibration, suggesting that stochasticity without structured uncertainty-aware computation is insufficient.
+Combining uncertainty-weighted attention with confidence-aware decision shaping yields the largest improvement among the evaluated components, consistent with their complementarity.
+To examine the design choice of where to apply uncertainty modulation within attention, we evaluate variants that gate the query (Q), key (K), value (V), or all three jointly (QKV).
+Fig.˜3 summarizes the calibration–accuracy trade-offs across SST-2, MNLI, and SQuAD 2.0.
+Overall, K-only achieves the lowest ECE on SST-2 and MNLI but can substantially reduce accuracy on SQuAD 2.0, suggesting overly aggressive evidence gating in span-style question answering.
+Q-only provides the most balanced behavior, improving calibration while largely preserving accuracy across tasks, whereas V-only best preserves, and in some cases slightly improves, accuracy while yielding more modest ECE gains.
+Joint QKV gating can further reduce ECE on some datasets but tends to over-regularize and degrade accuracy.
 6.6 Generalization Across BERT Models
@@
 Under matched evaluation protocols, UAT-Lite improves calibration across these pretrained backbones, yielding an average relative ECE reduction of approximately 32%.
-The largest gains occur for mid-sized and domain-specialized models, including BERT-base on SQuAD, BioBERT on PubMedQA, and ClinicalBERT on MedQA, where epistemic uncertainty arising from semantic ambiguity is prominent.
-Improvements are smaller for models with strong baseline calibration (e.g., RoBERTa-base on MNLI) and diminish or reverse for very small or highly overparameterized models.
-These results indicate that uncertainty-aware attention generalizes reliably within the BERT encoder family.
-Extending the mechanism to fundamentally different architectures, such as decoder-only or state-space models, remains an important direction for future work.
+The largest gains are observed for mid-sized and domain-specialized models, including BERT-base on SQuAD, BioBERT on PubMedQA, and ClinicalBERT on MedQA, where epistemic uncertainty associated with semantic ambiguity appears more prominent.
+By contrast, improvements are smaller for models with stronger baseline calibration (e.g., RoBERTa-base on MNLI) and diminish, or occasionally reverse, for very small or heavily overparameterized models.
+These results suggest that uncertainty-aware attention generalizes reliably within the BERT encoder family.
+Extending this mechanism to fundamentally different architectures, such as decoder-only or state-space models, remains an important direction for future work.
 Additional analyses.
-Additional analyses on adversarial robustness, linguistic uncertainty, clinical-domain transfer, and layer-wise uncertainty diagnostics are provided in the App.˜D.
+Additional analyses are provided in Apps.˜D, E and F, including threshold-based selective prediction, OOD evaluations (e.g., HANS), global and document-level MC dropout diagnostics, component ablations, penalty-form ablations, aggregation and normalization diagnostics, scale and geometry diagnostics for token uncertainty, negation diagnostics with qualitative attention illustration, behavioral analysis by linguistic bucket on MedQA, sensitivity analyses for and , character-level adversarial robustness, full linguistic uncertainty analysis, cross-model and scaling analyses, and implementation, complexity, latency, and reproducibility details.
 7 Conclusion
-We introduced UAT-Lite, an inference-time framework that integrates epistemic uncertainty into transformer computation through uncertainty-weighted attention and a layer-wise variance diagnostic.
-Across NLP benchmarks, UAT-Lite reduces Expected Calibration Error relative to unscaled fine-tuned BERT (e.g., average ECE 0.117 0.094) and improves selective prediction behavior, including higher coverage at fixed confidence thresholds and lower AURC.
-The method also exhibits stable calibration under distribution shift in MNLI matchedmismatched transfer.
-While inference-time Monte Carlo sampling incurs additional latency, the approach avoids the training and storage costs of ensemble-based alternatives and requires no modification of pretrained weights.
-Overall, UAT-Lite provides a complementary mechanism for uncertainty-aware transformer inference, bridging the gap between post-hoc calibration and Bayesian modeling for transformers.
+We introduced UAT-Lite, an inference-time framework that integrates epistemic uncertainty into transformer inference through uncertainty-weighted attention and a layer-wise variance diagnostic.
+Across general NLP benchmarks, UAT-Lite improves calibration relative to an unscaled fine-tuned BERT baseline and yields more reliable selective prediction behavior under fixed confidence thresholds.
+Under distribution shift (MNLI matchedmismatched and OOD evaluation suites), uncertainty-aware attention provides benefits that are complementary to output-level calibration: temperature scaling remains the strongest post-hoc baseline for ECE reduction, whereas UAT-Lite targets internal evidence aggregation and composes cleanly with TS when calibrated probabilities are required.
+Although Monte Carlo sampling increases inference latency, UAT-Lite requires neither retraining nor modification of pretrained weights, providing a practical bridge between lightweight post-hoc calibration methods and higher-cost Bayesian or ensemble-based alternatives.
 Limitations
-Our framework incurs additional computational cost due to Monte Carlo sampling, with a latency increase at .
-This overhead is appropriate for offline, batch, or document-level applications where confidence reliability is prioritized but is prohibitive for latency-critical settings.
-In practical deployments, this trade-off suggests that uncertainty-aware inference should be selectively enabled for high-risk or ambiguous inputs, rather than applied uniformly across all predictions.
-Accordingly, UAT-Lite is best viewed as an intermediate solution between lightweight post-hoc calibration and more expensive ensemble-based approaches, rather than a universal replacement.
-Finally, on clinical benchmarks (MedQA and PubMedQA), calibration gains are modest, reflecting dataset properties such as factual emphasis and limited ambiguity; these experiments are therefore treated as domain-transfer stress tests rather than clinical validation.
+Compute and latency.
+UAT-Lite introduces additional inference cost due to Monte Carlo dropout sampling.
+Under a single canonical end-to-end GPU benchmark (batch size , sequence length , sequential MC loop on one GPU), the overhead is substantial (e.g., approximately at ), whereas temperature scaling adds negligible cost (approximately ).
+As a result, UAT-Lite is not intended for latency-critical, always-on deployment.
+Instead, it is better suited to offline or batch settings, or to selectively triggered use (e.g., enabling uncertainty-aware attention only for low-confidence or high-risk inputs), where improved confidence reliability and internal evidence control may justify the additional runtime.
+Overall, UAT-Lite occupies an intermediate point between lightweight post-hoc calibration methods (e.g., TS) and higher-cost uncertainty methods (e.g., deep ensembles), trading additional compute for internal uncertainty-aware control signals and diagnostic capability.
+Compute-once token uncertainty (no per-layer recomputation).
+For efficiency and conceptual clarity, we compute token-level uncertainty once at the embedding stage (via stochastic embedding samples) and reuse this signal to modulate attention throughout the network; deeper layers are analyzed through variance attribution rather than by re-estimating after each mixing operation.
+Recomputing token uncertainty at intermediate layers is a potentially useful extension, but it would require defining token-level uncertainty after contextual mixing (e.g., residual connections and attention aggregation) and would further increase inference cost due to additional stochastic passes.
+We leave per-layer recomputation, and its associated cost–benefit trade-offs, to future work, especially in settings where additional compute is acceptable.
+Not a replacement for TS in-domain.
+Because temperature scaling is a strong lightweight calibrator for in-domain ECE, UAT-Lite should not be viewed as a universal substitute for TS.
+Our primary contribution is different: uncertainty-guided modulation of the internal evidence aggregation process through attention at inference time.
+Accordingly, we emphasize compositional use: UAT-Lite can be combined with TS (by fitting on development data using MC-mean logits) when the goal is to retain internal uncertainty-aware behavior while further improving output calibration.
+Clinical transfer is a stress test, not clinical validation.
+On clinical QA benchmarks (MedQA and PubMedQA), calibration gains can be modest, which may reflect dataset properties (e.g., a stronger emphasis on factual recall and less ambiguity in evidence aggregation) as well as evaluation constraints.
+We therefore treat these experiments as domain-transfer stress tests rather than evidence of clinical readiness.
+Any real-world clinical deployment would require task- and setting-specific validation, including broader coverage, robustness checks, and appropriate governance safeguards.
 Appendix
+This supplementary material provides additional experimental results, methodological details, and reproducibility information supporting the main paper.
+Appendix A presents the mathematical foundations of UAT-Lite, including the diagnostic layer-wise variance decomposition, its derivation, and the Bayesian interpretation of MC dropout.
+Appendix B provides the inference-time algorithm for uncertainty-aware attention.
+Appendix C details baselines, evaluation metrics (ECE, robustness metrics, selective prediction), and reporting conventions (seeds, fairness, confidence definitions).
+Appendix D reports additional experimental analyses:
+Appendix E evaluates generalization across BERT-family backbones and model sizes, including cross-domain calibration and scaling effects.
+Appendix F documents implementation details, compute/latency overhead, hardware setup, and code availability for reproducibility.
+Appendix G provides an extended related work discussion.
+Appendix H provides an AI usage statement.
 Appendix A Mathematical Details
-In this section, we provide the formal mathematical foundations underlying the uncertainty modeling and diagnostic analyses used in UAT-Lite.
+In this section, we present the formal mathematical foundations underlying the uncertainty modeling and diagnostic analyses used in UAT-Lite.
 We first introduce a layer-wise variance decomposition that attributes predictive uncertainty across transformer depth under stochastic inference.
 We then provide a proof of the decomposition, clarify its assumptions, and relate Monte Carlo dropout to approximate Bayesian inference.
-Finally, we discuss practical and interpretive implications of layer-wise uncertainty attribution for calibration, intervention, and uncertainty-aware inference.
-These analyses are intended as diagnostic and explanatory tools that complement the empirical results in the main paper, rather than as claims of exact posterior inference.
+Finally, we discuss the practical and interpretive implications of layer-wise uncertainty attribution for calibration, intervention, and uncertainty-aware inference.
+These analyses serve as diagnostic and explanatory tools that complement the empirical results in the main paper, rather than as claims of exact posterior inference.
 A.1 Layer-Wise Uncertainty Decomposition
-To better understand how predictive uncertainty emerges within the transformer, we analyze its distribution across layers under stochastic inference.
+To better understand how predictive uncertainty arises within the transformer, we analyze its distribution across layers under stochastic inference.
 Normalized variance contributions.
@@
 such that contributions across layers sum to . All percentages reported below correspond to these normalized values.
-Table 5 presents representative layer-wise uncertainty decompositions for answerable and unanswerable SQuAD examples, illustrating how predictive uncertainty is distributed across transformer depth rather than manifesting solely at the output layer.
-We emphasize that these examples are illustrative diagnostics rather than statistical summaries, and are intended to provide qualitative insight into uncertainty propagation rather than population-level guarantees.
+Table 3 presents representative layer-wise uncertainty decompositions for answerable and unanswerable SQuAD examples, illustrating how predictive uncertainty is distributed across transformer depth rather than appearing only at the output layer.
+We emphasize that these examples are illustrative diagnostics rather than statistical summaries and are intended to provide qualitative insight into uncertainty propagation rather than population-level guarantees.
 For the unanswerable example, uncertainty grows progressively through the self-attention stack.
@@
 This pattern is consistent with uncertainty emerging during later reasoning stages when sufficient evidence is unavailable, rather than arising only at the final decision layer.
-In contrast, the answerable example exhibits a similar normalized uncertainty distribution but with a different internal structure.
+In contrast, the answerable example exhibits a similar normalized uncertainty distribution but a different internal structure.
 Mid-layer attention contributes more strongly to evidence aggregation (8.7%), while late-layer uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
@@
 | Total | 1.000 | 100% | Confident prediction | |
-A.2 Proof of Theorem 1
+A.2 Derivation of Proposition 1
 Proof.
-We derive the layer-wise variance decomposition by applying the law of total variance recursively across transformer layers.
+We derive a diagnostic layer-wise variance decomposition by recursively applying the law of total variance across transformer depth.
+The purpose of this derivation is interpretive: it provides a layer-indexed summary of where predictive variance arises or amplifies during stochastic inference.
 Setup and notation.
-Let denote the model’s final prediction (e.g., class probability or logit), and let denote the hidden representation at layer corresponding to the [CLS] token.
-Stochasticity is introduced via Monte Carlo dropout at inference time.
+Let denote the model’s final prediction (e.g., a class probability or logit), and let denote the hidden representation at layer corresponding to the [CLS] token.
+Stochasticity is introduced through Monte Carlo dropout at inference time.
 Each transformer layer implements a stochastic mapping
-where denotes dropout-induced noise at layer . We assume finite second moments for all noise distributions but do not assume Gaussianity or independence across layers.
-Layer-wise variance contribution.
-We define the variance contribution of layer as
-Recursive variance decomposition.
-Applying the law of total variance at the final layer yields
-Applying the same decomposition recursively yields an additive expansion over layers:
+where denotes dropout-induced noise at layer . We assume finite second moments so that all conditional variances below are well defined.
+Recursive application of total variance.
+Applying the law of total variance at the final layer gives
+The same identity can then be applied recursively to the second term by conditioning on progressively earlier hidden representations .
+This yields a recursive decomposition of predictive variance across depth.
+Diagnostic layer-wise contributions.
+Motivated by this recursion, we define the layer-wise diagnostic contribution of layer as
+These quantities summarize the predictive variability associated with the stochasticity introduced at each layer, conditioned on the representation propagated from preceding layers.
 Interpretation.
-This decomposition provides an attribution of predictive uncertainty across layers under conditional expectations.
-While inter-layer dependencies may introduce higher-order interaction effects, the decomposition remains useful as an interpretable diagnostic of where uncertainty is amplified during inference.
+Under this construction, predictive variance admits a recursive layer-indexed decomposition, and the normalized values
+provide an interpretable summary of how uncertainty is distributed across depth.
+We emphasize that these contributions are intended as diagnostic attributions: they are useful for identifying where uncertainty concentrates or amplifies during inference, but they should not be interpreted as unique causal effects or as a novel variance identity.
 ∎
@@
 Uncertainty interpretation.
-The variability across predictions reflects epistemic uncertainty arising from model and data limitations, which is reducible with additional information.
+Variability across predictions reflects epistemic uncertainty arising from model and data limitations and is, in principle, reducible with additional information.
 Relevance to UAT-Lite.
@@
 As shown in Table 3, unanswerable SQuAD examples concentrate substantial uncertainty in late layers (9–11), consistent with ambiguity emerging during final reasoning stages.
-Answerable examples exhibit comparable late-layer contributions, but with uncertainty reflecting confident task-level reasoning rather than unresolved ambiguity.
+Answerable examples exhibit comparable late-layer contributions, but the uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
 Intervention.
-Layer-wise attribution suggests that calibration strategies could in principle target high-variance components.
-Our ablation study (Fig.˜4) shows that uncertainty introduced at the attention level yields a 21.1% improvement in calibration, while embedding-level uncertainty contributes negatively (-3.1%), supporting selective intervention rather than uniform stochasticity.
+Layer-wise attribution suggests that calibration strategies could, in principle, target high-variance components.
+Our ablation study (Fig.˜5) shows that uncertainty introduced at the attention level yields a 21.1% improvement in calibration, while embedding-level uncertainty contributes negatively (-3.1%), supporting selective intervention rather than uniform stochasticity.
 Interpretability.
@@
 For complex syntactic constructions, uncertainty is elevated in early and mid layers due to structural processing, while late layers consolidate the final decision.
+Scope and limitation (per-layer recomputation).
+While we analyze where uncertainty concentrates across depth via layer-wise variance attribution (Table 3), our UAT-LITE intervention computes token uncertainty once (from stochastic forward passes) and reuses it across layers for efficiency.
+Recomputing token uncertainty after each layer would require defining uncertainty over mixed representations after contextualization and would increase inference cost roughly in proportion to the number of layers; we leave such layer-wise recomputation to future work.
 A.5 Practical Inference-Time Procedure
@@
 The mean is used as the final prediction, while the variance quantifies epistemic uncertainty.
-Uncertainty estimates influence inference by modulating attention behavior under representation ambiguity and enabling confidence-aware decision shaping at the output level.
+Uncertainty estimates influence inference by modulating attention under representation ambiguity and enabling confidence-aware decision shaping at the output level.
 All uncertainty mechanisms operate strictly at inference time and do not modify training dynamics or learned parameters.
@@
 Appendix C Baselines and Metrics
-We compare against BERT-base (110M parameters), MC Dropout (, ), Temperature Scaling (TS), and Deep Ensembles (5 independently fine-tuned models).
-Temperature scaling is treated as a strong post-hoc calibration baseline that operates purely at the output level and does not modify internal model computation.
-Because TS rescales logits without affecting representations, it is compatible with uncertainty-aware inference and could in principle be applied on top of UAT-Lite.
-In this work, however, we report TS as a standalone baseline and do not include a stacked (UAT-Lite + TS) variant in the main tables, in order to isolate the effects of uncertainty-aware internal computation from output-level rescaling.
+We compare against BERT-base (110M parameters), MC Dropout (, ), TS, and Deep Ensembles (5 independently fine-tuned models).
+TS is treated as a strong post-hoc calibration baseline that operates purely at the output level and does not modify internal model computation.
+Because TS rescales logits without affecting representations, it is compatible with uncertainty-aware inference and can be applied on top of UAT-Lite.
+Accordingly, we report both TS as a standalone baseline and the stacked variant UAT-Lite +TS to assess composability.
+For UAT-Lite +TS, the temperature is fit on the development set using the MC-mean logits produced by UAT-Lite and applied at test time.
+To isolate the effect of uncertainty-aware internal computation, we compare UAT-Lite against non-stacked baselines (e.g., Base, MC Dropout, and TS), and treat UAT-Lite +TS as an explicit compositional setting rather than the core mechanism itself.
 All baselines use the same pretrained backbone, fine-tuning recipe, and evaluation code unless otherwise stated.
@@
 where predictions are partitioned into fixed-width confidence bins over .
-For each bin, denotes empirical accuracy and denotes the mean predicted confidence.
-Here denotes the number of confidence bins used for calibration evaluation, distinct from the number of Monte Carlo samples used for uncertainty estimation.
+For each bin, denotes empirical accuracy, and denotes mean predicted confidence.
+Here, denotes the number of confidence bins used for calibration evaluation, which is distinct from the number of Monte Carlo samples used for uncertainty estimation.
 Lower ECE indicates better calibration.
@@
 SQuAD 2.0 evaluation protocol.
-We evaluate SQuAD 2.0 on the answerability prediction task, where the model must determine whether a question is answerable given the context.
-This binary classification task is well-suited for calibration analysis.
+We evaluate SQuAD 2.0 on the answerability prediction task, in which the model determines whether a question is answerable given the context.
+This binary classification setting is well suited to calibration analysis.
 We report answerability prediction accuracy alongside ECE in Table 1.
-Note that span-level extraction (predicting answer start/end positions) is a separate task and orthogonal to our uncertainty calibration contribution.
-Our focus is on calibrated confidence scores for the answerability decision, not optimizing span extraction performance.
+Span-level extraction (predicting answer start and end positions) is a separate task and is orthogonal to our uncertainty calibration contribution.
+Our focus is on calibrated confidence scores for the answerability decision rather than on optimizing span extraction performance.
 Distribution shift robustness metrics.
@@
 First, we measure calibration drift:
-where smaller (or negative) values indicate improved robustness under shift. Second, we report a calibration robustness score defined as the average ECE across ID and OOD settings:
+where smaller (or negative) values indicate greater robustness under shift. Second, we report a calibration robustness score defined as the average ECE across ID and OOD settings:
 with lower values indicating more stable calibration across domains.
@@
 We additionally report selective coverage at fixed confidence thresholds and the Area Under the Risk–Coverage Curve (AURC; lower is better).
-Confidence thresholds are selected on the validation set and held fixed at test time for all methods.
+Confidence thresholds are selected on the validation set for each method and then held fixed at test time.
+We also report admitted counts () at fixed thresholds to make score-rescaling effects (e.g., under TS) explicit.
 MC Dropout baseline configuration.
 The MC Dropout baseline uses a uniform dropout rate of across all layers, consistent with prior calibration studies.
-While more aggressive or component-specific dropout tuning could further improve MC Dropout performance, such tuning increases implementation complexity and reduces comparability with standard baselines.
+Although more aggressive or component-specific dropout tuning could further improve MC Dropout performance, such tuning increases implementation complexity and reduces comparability with standard baselines.
 In contrast, UAT-Lite explicitly leverages structured uncertainty within the attention mechanism rather than relying solely on increased stochasticity.
@@
 Component-specific MC Dropout control.
-To explicitly disentangle the effect of structured stochasticity from uncertainty-aware attention, we additionally evaluate an MC Dropout baseline using the same component-specific dropout rates as UAT-Lite (0.1 embeddings / 0.2 attention / 0.3 feed-forward), but without attention modulation.
+To disentangle the effect of structured stochasticity from uncertainty-aware attention, we additionally evaluate an MC Dropout baseline using the same component-specific dropout rates as UAT-Lite (0.1 embeddings / 0.2 attention / 0.3 feed-forward), but without attention modulation.
 As shown in Table 1, this control does not improve calibration relative to uniform MC Dropout or the BERT baseline, indicating that the gains of UAT-Lite arise from uncertainty-weighted attention rather than tuned dropout alone.
@@
 ECE is computed with fixed-width bins for all methods.
-Temperature scaling (and UAT-Lite +TS) fits a single scalar temperature on the validation set only, and is applied at test time without any access to test labels.
-OOD calibration metrics (e.g., ECE and Robustness) are computed using the identical pipeline on ID and OOD sets to avoid binning or confidence-definition mismatch.
+Temperature scaling (and UAT-Lite +TS) fits a single scalar temperature on the validation set only and is applied at test time without access to test labels.
+OOD calibration metrics (e.g., ECE and Robustness) are computed using the same pipeline on ID and OOD sets to avoid binning or confidence-definition mismatch.
 Random seeds and variability.
 Main-paper results (general calibration, distribution-shift robustness, and selective prediction) are reported as mean standard deviation over five random seeds.
-Appendix diagnostic analyses—including qualitative attention visualizations, representative layer-wise uncertainty decompositions, and linguistic or adversarial case studies—are reported using a single seed to reduce computational cost and are explicitly labeled as single-seed in the corresponding captions.
+Appendix diagnostic analyses, including qualitative attention visualizations, representative layer-wise uncertainty decompositions, and linguistic or adversarial case studies, are reported using a single seed to reduce computational cost and are explicitly labeled as single-seed in the corresponding captions.
 Deterministic baselines exhibit negligible variance because inference is fully deterministic when evaluated from a fixed fine-tuned checkpoint reused across seeds.
 Appendix D Additional Experimental Results
-This section presents additional analyses that complement the main experimental results by providing deeper insight into the behavior of UAT-Lite in specialized and fine-grained settings.
-Specifically, we examine calibration performance in clinical benchmarks, where decision reliability is critical, and conduct a detailed linguistic analysis to assess how uncertainty-aware attention responds to diverse sources of ambiguity, structure, and semantic inconsistency.
-Together, these experiments offer a more nuanced understanding of the strengths and limitations of the proposed framework beyond aggregate calibration metrics.
-D.1 Sensitivity Analysis of and
-We examine the sensitivity of UAT-Lite to its two primary inference-time hyperparameters: the uncertainty weighting coefficient and the number of Monte Carlo samples .
-A small, representative grid is evaluated with and on SQuAD 2.0, which is particularly sensitive to epistemic uncertainty due to the presence of both answerable and unanswerable questions.
+This section presents additional analyses that complement the main experimental results by providing finer-grained insight into the behavior of UAT-Lite.
+Specifically, we examine selective prediction under fixed thresholds, calibration behavior under stronger OOD shifts, and settings that reveal how uncertainty-aware attention responds to ambiguity, structure, and semantic inconsistency.
+Together, these analyses provide a more nuanced view of the strengths and limitations of the proposed framework beyond aggregate calibration metrics.
+D.1 Additional Threshold-Based Selective Prediction Results
+In addition to the broader OOD suites reported in the main text, we provide threshold-based selective prediction results on MNLI (N=4,908) and SQuAD 2.0 (N=5,937) using the same evaluation protocol.
+Confidence is defined as , and for thresholds we report coverage (Cov@), admitted accuracy (Acc@), admitted counts (), and AURC.
+Table 4 serves two purposes.
+First, it shows that the selective prediction behavior observed under stronger OOD conditions also appears in standard evaluation settings.
+Second, it clarifies how post-hoc temperature scaling (TS) changes the admitted fraction at fixed thresholds.
+In particular, TS sharply compresses high-confidence regions (e.g., Cov@0.9 on SQuAD), while maintaining high admitted accuracy.
+This behavior reflects output probability rescaling rather than structural changes in evidence aggregation.
+By contrast, UAT-LITE modifies confidence through uncertainty-aware attention modulation.
+As a result, changes in coverage reflect altered internal evidence weighting rather than purely scalar logit rescaling.
+These results complement the OOD analysis in the main text by showing the same protocol under milder distribution conditions.
+Dataset Method AURC Cov Acc Cov Acc Cov Acc MNLI Baseline 0.3857 0.8822 1893 0.5591 0.8455 2744 0.6999 0.8032 3435 0.1446 Base+TS 0.1795 0.9432 881 0.3820 0.8853 1875 0.5515 0.8493 2707 0.1441 MC Dropout (M=10) 0.3048 0.8984 1496 0.4890 0.8596 2400 0.6447 0.8186 3164 0.1478 UAT-LITE (M=10) 0.3195 0.8929 1568 0.5026 0.8541 2467 0.6524 0.8154 3202 0.1483 UAT-LITE+TS 0.1746 0.9428 857 0.3714 0.8826 1823 0.5383 0.8471 2642 0.1479 SQuAD 2.0 Baseline 0.4745 0.7000 2817 0.6468 0.6974 3840 0.7780 0.6850 4619 0.3186 Base+TS 0.0000 0.0000 0 0.0372 0.6787 221 0.1816 0.6698 1078 0.3186 MC Dropout (M=10) 0.3901 0.7003 2316 0.5875 0.6995 3488 0.7349 0.6890 4363 0.3150 UAT-LITE (M=10) 0.4051 0.6998 2405 0.5979 0.7000 3550 0.7408 0.6885 4398 0.3177 UAT-LITE+TS 0.0000 0.0000 0 0.0487 0.6574 289 0.1728 0.6676 1026 0.3177
+D.2 OOD Threshold-Based Selective Prediction (HANS/ANLI/SNLI/IMDb)
+To further evaluate uncertainty quality under stronger distribution shifts, we conduct threshold-based selective prediction experiments across four widely used out-of-distribution (OOD) evaluation suites: HANS (), ANLI (), SNLI (), and IMDb ().
+The full results are reported in Table 5.
+Selective prediction evaluates whether a model can abstain on low-confidence examples while maintaining high accuracy on the retained subset.
+For a confidence threshold , we admit predictions only when the model confidence exceeds .
+For each threshold, we report coverage (Cov@), admitted accuracy (Acc@), admitted count (), and AURC.
+Table 5 consolidates results across all OOD suites in a single view, with Panels A–D corresponding to HANS, ANLI, SNLI, and IMDb, respectively.
+HANS and ANLI represent particularly challenging shifts designed to break common NLI heuristics, whereas SNLI and IMDb provide additional cross-domain evaluation settings.
+Across datasets, the table shows how different uncertainty estimation methods affect selective prediction behavior.
+In particular, TS often changes probability calibration without altering the predicted label, which can substantially shift coverage at fixed thresholds.
+Accordingly, differences in Cov@ should be interpreted as changes in the confidence distribution rather than in classification decisions.
+This appendix therefore complements the MNLI transfer analysis by providing a broader view of uncertainty-guided filtering under diverse OOD conditions.
+| Method | Acc | ECE | NLL | Cov@.9 | Acc@.9 | N@.9 | Cov@.8 | Acc@.8 | N@.8 | Cov@.7 | Acc@.7 | N@.7 | AURC |
+| Panel A: HANS OOD (N=30,000) | |||||||||||||
+| Baseline | 0.4808 | 0.3092 | 1.0581 | 0.2209 | 0.4831 | 6626 | 0.5427 | 0.4930 | 16281 | 0.7502 | 0.4840 | 22507 | 0.5287 |
+| Base+TS | 0.4836 | 0.2283 | 0.8737 | 0.0625 | 0.3952 | 1875 | 0.2333 | 0.4856 | 6999 | 0.5415 | 0.4942 | 16246 | 0.5268 |
+| MC Dropout | 0.4810 | 0.2506 | 0.9256 | 0.0883 | 0.4392 | 2650 | 0.3428 | 0.4816 | 10284 | 0.5929 | 0.4799 | 17788 | 0.5322 |
+| UAT-LITE | 0.4807 | 0.2395 | 0.9074 | 0.0687 | 0.4469 | 2061 | 0.3039 | 0.4650 | 9117 | 0.5583 | 0.4646 | 16749 | 0.5385 |
+| UAT-LITE+TS | 0.4817 | 0.1908 | 0.8222 | 0.0210 | 0.3386 | 629 | 0.1312 | 0.4554 | 3935 | 0.3951 | 0.4673 | 11853 | 0.5386 |
+| Panel B: ANLI OOD (N=3,200) | |||||||||||||
+| Baseline | 0.3109 | 0.4209 | 1.9074 | 0.2050 | 0.2668 | 656 | 0.3978 | 0.2875 | 1273 | 0.5747 | 0.2953 | 1839 | 0.7130 |
+| Base+TS | 0.3109 | 0.3413 | 1.5450 | 0.0541 | 0.2370 | 173 | 0.2013 | 0.2609 | 644 | 0.3769 | 0.2886 | 1206 | 0.7126 |
+| MC Dropout | 0.3106 | 0.3916 | 1.7681 | 0.1456 | 0.2446 | 466 | 0.3231 | 0.2834 | 1034 | 0.5094 | 0.2914 | 1630 | 0.7141 |
+| UAT-LITE | 0.3091 | 0.4040 | 1.8222 | 0.1778 | 0.2566 | 569 | 0.3516 | 0.2836 | 1125 | 0.5319 | 0.2955 | 1702 | 0.7130 |
+| UAT-LITE+TS | 0.3091 | 0.3443 | 1.5591 | 0.0656 | 0.2381 | 210 | 0.2112 | 0.2692 | 676 | 0.3847 | 0.2851 | 1231 | 0.7124 |
+| Panel C: SNLI OOD (N=9,824) | |||||||||||||
+| Baseline | 0.6810 | 0.0684 | 0.7757 | 0.2579 | 0.9013 | 2534 | 0.4540 | 0.8377 | 4460 | 0.6162 | 0.7907 | 6054 | 0.1772 |
+| Base+TS | 0.6810 | 0.0179 | 0.7497 | 0.0618 | 0.9176 | 607 | 0.2513 | 0.9032 | 2469 | 0.4379 | 0.8440 | 4302 | 0.1760 |
+| MC Dropout | 0.6819 | 0.0328 | 0.7497 | 0.1766 | 0.9182 | 1735 | 0.3727 | 0.8651 | 3661 | 0.5408 | 0.8095 | 5313 | 0.1752 |
+| UAT-LITE | 0.6759 | 0.0474 | 0.7650 | 0.1995 | 0.9122 | 1960 | 0.3954 | 0.8540 | 3884 | 0.5610 | 0.8044 | 5511 | 0.1792 |
+| UAT-LITE+TS | 0.6759 | 0.0195 | 0.7534 | 0.0693 | 0.9383 | 681 | 0.2438 | 0.8944 | 2395 | 0.4274 | 0.8445 | 4199 | 0.1783 |
+| Panel D: IMDb OOD (N=25,000) | |||||||||||||
+| Baseline | 0.8281 | 0.1000 | 0.4904 | 0.7780 | 0.8925 | 19451 | 0.8641 | 0.8687 | 21602 | 0.9173 | 0.8538 | 22932 | 0.0640 |
+| Base+TS | 0.8281 | 0.0442 | 0.3959 | 0.5960 | 0.9325 | 14899 | 0.7683 | 0.8950 | 19208 | 0.8634 | 0.8688 | 21586 | 0.0640 |
+| MC Dropout | 0.8195 | 0.0957 | 0.4845 | 0.7367 | 0.8938 | 18417 | 0.8405 | 0.8673 | 21012 | 0.9014 | 0.8495 | 22536 | 0.0695 |
+| UAT-LITE | 0.8186 | 0.1054 | 0.5078 | 0.7674 | 0.8862 | 19184 | 0.8570 | 0.8626 | 21426 | 0.9109 | 0.8463 | 22772 | 0.0694 |
+| UAT-LITE+TS | 0.8186 | 0.0501 | 0.4117 | 0.5781 | 0.9279 | 14453 | 0.7598 | 0.8882 | 18995 | 0.8583 | 0.8621 | 21457 | 0.0694 |
+D.3 Global Document-Level MC Dropout Diagnostics
+To contextualize the behavior of uncertainty-weighted attention, we evaluate standard global/document-level MC Dropout applied to the same BERT-base checkpoint used in the main results.
+Here, dropout is injected at inference time without modifying the attention mechanism, and predictive uncertainty is estimated from Monte Carlo sampling ().
+Fig.˜4 summarizes calibration (ECE, Brier, NLL), selective prediction (AURC, Acc@70/80/90), and accuracy across SQuAD 2.0, MNLI, and SST-2 (absolute values shown in each cell; color denotes row-normalized relative performance).
+The purpose of this analysis is diagnostic rather than competitive.
+Global MC Dropout improves over the unscaled baseline in calibration (e.g., ECE decreases across datasets), but it remains materially behind temperature scaling in marginal calibration (see the TS gaps reported elsewhere), confirming that naive stochastic inference alone does not recover TS-level calibration.
+Importantly, this comparison isolates the effect of where uncertainty is injected.
+Unlike global MC Dropout which affects output variance while leaving evidence aggregation unchanged, UAT-LITE incorporates uncertainty directly into the attention logits, modifying evidence aggregation rather than only output variability.
+The diagnostic therefore clarifies that improvements observed in the main text are not attributable merely to Monte Carlo sampling, but to attention-level modulation.
+D.4 Heuristic Sensitivity Analysis on HANS
+To localize failure modes under distribution shift, we provide a per-heuristic breakdown on the HANS benchmark.
+HANS isolates three common heuristic shortcuts: lexical overlap, subsequence, and constituent structure.
+Table 6 reports overall accuracy by heuristic category (top block) and non-entailment-only accuracy (bottom block), which is the more challenging and shortcut-prone regime.
+The top block shows that overall per-heuristic differences are relatively modest across methods.
+However, the non-entailment-only breakdown is more revealing, because this subset isolates cases where shallow heuristics often produce overconfident errors.
+In this regime, UAT-LITE and especially UAT-LITE+TS show clear gains over the baseline, with the largest improvements on constituent-based examples.
+These results suggest that uncertainty-aware attention is most useful not in easy heuristic-aligned cases, but in examples where shortcut-driven evidence is misleading and the model must suppress superficially plausible matches.
+Overall, this diagnostic provides a more localized view of robustness under shift than aggregate HANS accuracy alone.
+| Method | Lex. Overlap | Subsequence | Constituent |
+| Baseline | 0.5043 | 0.4632 | 0.4750 |
+| Base+TS | 0.5104 | 0.4622 | 0.4783 |
+| MC Dropout | 0.5000 | 0.4642 | 0.4787 |
+| UAT-LITE | 0.4835 | 0.4777 | 0.4810 |
+| UAT-LITE+TS | 0.4859 | 0.4710 | 0.4881 |
+| Non-entailment only | |||
+| Baseline | 0.1804 | 0.2358 | 0.3428 |
+| Base+TS | 0.2230 | 0.2714 | 0.3824 |
+| MC Dropout | 0.2132 | 0.2408 | 0.4158 |
+| UAT-LITE | 0.2960 | 0.2722 | 0.5580 |
+| UAT-LITE+TS | 0.3548 | 0.3184 | 0.6100 |
+D.5 Component Ablation of Uncertainty Injection
+Fig.
+5 evaluates the contribution of individual components of UAT-LITE on SQuAD 2.0 (validation).
+We isolate: (i) baseline inference, (ii) MC sampling without attention modulation, (iii) uncertainty-weighted attention, and (iv) the full combined model.
+The primary metric is Expected Calibration Error (ECE), with accuracy reported for reference.
+Two observations are noteworthy.
+First, uncertainty-weighted attention produces the largest reduction in ECE relative to baseline, indicating that incorporating epistemic variance into the attention logits directly improves calibration.
+Second, combining sampling with attention modulation yields further improvements, suggesting that stochastic sampling and structural uncertainty injection are complementary.
+Notably, accuracy remains stable across variants, indicating that improvements in calibration are not driven by trivial accuracy trade-offs.
+This ablation confirms that gains observed in the main text are not attributable to Monte Carlo sampling alone, but specifically to attention-level uncertainty weighting.
+D.6 Penalty Functional Form Ablation
+To assess whether the exponential uncertainty penalty is essential, we compare three modulation forms under identical settings (, , SQuAD 2.0, single seed): (i) exponential: , (ii) linear (clipped): , and (iii) reciprocal: .
+Fig.
+6 reports the results.
+The three forms yield broadly comparable accuracy and calibration behavior, indicating that the overall effect stems from uncertainty-aware modulation rather than a single functional choice.
+We retain the exponential form due to its smoothness, guaranteed positivity (without clipping), and interpretation as an additive logit penalty (), which offers a convenient and numerically well-behaved mechanism for downweighting high-uncertainty evidence.
+D.7 Aggregation and Normalization Diagnostics
+This section probes the robustness of two implementation choices underlying UAT-Lite: (i) how we aggregate Monte Carlo variance across draws, and (ii) where normalization is applied prior to uncertainty measurement.
+We evaluate these choices using the MNLI negation diagnostic, which perturbs inputs without changing the original labels, and we track both label-dependent signals (error change) and label-free stability signals (confidence change) to ensure that observed behaviors are not artifacts of a particular aggregation heuristic or a specific LayerNorm placement.
+Aggregation-statistic sanity check (no LayerNorm).
+Table 7 (top block) compares three variance aggregation statistics (mean-std, RMS-std, median-std) on the MNLI negation diagnostic (, seed, ).
+We report changes in confidence and error (dConf/dErr), computed as (negated original), along with ECE on the sampled subset (SmpECE) and development ECE (DevECE).
+Across aggregation schemes, dConf and dErr are tightly clustered, indicating that the negation sensitivity signal is stable to the aggregation choice.
+Although SmpECE/DevECE vary slightly across statistics, the qualitative conclusions and effect directions remain unchanged.
+LayerNorm invariance.
+Table 7 (bottom block) tests whether applying LayerNorm before uncertainty measurement changes the diagnostic outcomes.
+For each aggregation statistic, we compare dConf, dErr, and DevECE with and without LayerNorm (“+LN”).
+We observe nearly unchanged dConf and comparable dErr across normalization conditions, while DevECE improves consistently under +LN.
+Overall, these results suggest that UAT-Lite is not sensitive to the precise LayerNorm placement and that the reported negation behaviors are not normalization artifacts.
+Variant AccOrig AccNeg dConf dErr SmpECE DevECE Aggregation-statistic sanity check (no LayerNorm) Base Det.
+0.7050 0.4950 0.1313 – mean-std 0.7000 0.4900 0.1123 0.0595 RMS-std 0.6800 0.4750 0.1199 0.0576 median-std 0.6850 0.5000 0.1056 0.0584 LayerNorm-invariance check (with/without LN) Variant dConf dConf+LN dErr dErr+LN DevECE DevECE+LN mean-std 0.0595 0.0466 RMS-std 0.0576 0.0455 median-std 0.0584 0.0482
+D.8 Scale and Geometry Diagnostics for Token Uncertainty
+This appendix section examines whether token-level uncertainty computed from embedding-space stochasticity reflects a meaningful epistemic signal or could be explained by simple geometric artifacts (e.g., global embedding rescaling, anisotropy, or magnitude effects).
+To stress-test this possibility, we conduct three complementary diagnostics on representative tasks (MNLI and SQuAD) under the same inference-time configuration used in the main experiments.
+First, we test robustness under explicit embedding rescaling () and quantify the resulting changes in task metrics and token-uncertainty rankings.
+Second, we evaluate alignment between token uncertainty and output-space epistemic quantities derived from MC dropout (predictive entropy and mutual information), while controlling for embedding norm to rule out trivial magnitude explanations.
+Third, we ablate alternative uncertainty definitions to verify that conclusions do not depend on a single aggregation statistic.
+Together, these analyses clarify what aspects of are robust, what is scale-sensitive, and why the observed behavior is not reducible to embedding magnitude alone.
+D.8.1 Embedding-Rescaling Invariance Test
+This diagnostic tests whether token uncertainty derived from embedding-space stochasticity could be explained by simple embedding-scale effects.
+We rescale the input token embeddings at inference time by a constant factor while keeping all other settings fixed, and rerun evaluation on MNLI and SQuAD.
+Table 8 reports task-level metrics and deltas relative to the unscaled baseline (), along with (i) token-uncertainty rank stability via Spearman correlation and (ii) attention-pattern stability via cosine similarity between the resulting attention weights.
+Two observations are salient.
+First, attention modulation remains highly stable under rescaling (Attn CosSim across all ), indicating that the induced attention patterns are largely preserved even under substantial changes in embedding magnitude.
+Second, token-uncertainty rankings are not strictly scale-invariant (– for ), and calibration metrics can shift under extreme rescaling (e.g., MNLI ECE up to at , while SQuAD ECE improves at ).
+Thus, while the attention behavior is robust, the uncertainty signal exhibits bounded scale sensitivity, motivating the complementary sanity checks and uncertainty-definition ablations below.
+D.8.2 Epistemic Alignment via Correlation Checks
+To test whether token-level uncertainty aligns with more standard output-space epistemic indicators, we compute (under the same MC dropout samples) predictive entropy and mutual information , and measure their association with .
+Table 9 reports both Pearson and Spearman correlations.
+In addition, to rule out the possibility that is simply tracking embedding magnitude, we report correlations with embedding norm and partial correlations between and controlling for .
+On MNLI, shows consistent monotonic association with output-space uncertainty (Spearman with entropy and with MI), while linear correlations are small but statistically reliable given the large number of tokens.
+On SQuAD, correlations are weaker, suggesting task-dependent alignment between token- and output-space uncertainty.
+Importantly, is essentially uncorrelated with embedding norm on both datasets (Pearson near ; Spearman near ), and the norm-controlled partial correlations remain nearly unchanged, indicating that the signal is not explained by embedding magnitude alone.
+D.8.3 Alternative Uncertainty Definitions: Robustness Ablation
+Finally, we evaluate whether the results depend on a particular choice of uncertainty definition.
+Beyond the default mean-std, we compare alternatives that are commonly used to summarize multivariate dispersion under stochastic embeddings: cov-trace (trace of the covariance), cv (coefficient of variation), and norm-cv (a normalized CV variant).
+Table 10 shows that performance and calibration remain stable across these variants on both MNLI and SQuAD (differences are small and do not change the qualitative conclusions), suggesting that UAT-LITE does not hinge on a single arbitrary aggregation of embedding-space variability.
+| Dataset | ECE | NLL | Acc | AURC | ECE | NLL | Acc | |
+| MNLI | ||||||||
+| MNLI | 0.5 | 0.0117 | 0.5335 | 0.7773 | 0.0836 | -0.0048 | +0.0138 | -0.0100 |
+| MNLI | 1.0 | 0.0165 | 0.5197 | 0.7873 | 0.0783 | +0.0000 | +0.0000 | +0.0000 |
+| MNLI | 2.0 | 0.0260 | 0.5418 | 0.7753 | 0.0864 | +0.0095 | +0.0221 | -0.0120 |
+| MNLI | 4.0 | 0.0212 | 0.5553 | 0.7613 | 0.0924 | +0.0047 | +0.0356 | -0.0260 |
+| SQuAD | ||||||||
+| SQuAD | 0.5 | 0.3216 | 1.3102 | 0.5075 | 0.4923 | -0.0123 | -0.0537 | +0.0035 |
+| SQuAD | 1.0 | 0.3339 | 1.3639 | 0.5040 | 0.4913 | +0.0000 | +0.0000 | +0.0000 |
+| SQuAD | 2.0 | 0.3250 | 1.2799 | 0.5020 | 0.4944 | -0.0089 | -0.0840 | -0.0020 |
+| SQuAD | 4.0 | 0.2944 | 1.1441 | 0.5060 | 0.4937 | -0.0395 | -0.2198 | +0.0020 |
+| Dataset | Spearman | Attn CosSim | |
+| MNLI | |||
+| MNLI | 0.5 | 0.2332 | 0.9869 |
+| MNLI | 2.0 | 0.2120 | 0.9842 |
+| MNLI | 4.0 | 0.2015 | 0.9792 |
+| SQuAD | |||
+| SQuAD | 0.5 | 0.1522 | 0.9922 |
+| SQuAD | 2.0 | 0.0998 | 0.9913 |
+| SQuAD | 4.0 | 0.1245 | 0.9851 |
+| Dataset | – (P) | – (S) | –MI (P) | –MI (S) | – (P) | – (S) | ||
+| MNLI | 0.0580 | 0.3677 | 0.0759 | 0.4137 | -0.0071 | -0.0236 | 0.0583 | 0.0759 |
+| SQuAD | -0.0163 | -0.0222 | 0.0055 | 0.1694 | -0.0256 | -0.0510 | -0.0157 | 0.0041 |
+| Dataset | Variant | ECE | NLL | Acc | AURC |
+| MNLI | |||||
+| MNLI | mean-std | 0.0158 | 0.5128 | 0.7908 | 0.0775 |
+| MNLI | cov-trace | 0.0158 | 0.5128 | 0.7908 | 0.0775 |
+| MNLI | cv | 0.0165 | 0.5197 | 0.7873 | 0.0783 |
+| MNLI | norm-cv | 0.0158 | 0.5128 | 0.7908 | 0.0775 |
+| SQuAD | |||||
+| SQuAD | mean-std | 0.3377 | 1.3866 | 0.5065 | 0.4940 |
+| SQuAD | cov-trace | 0.3377 | 1.3866 | 0.5065 | 0.4940 |
+| SQuAD | cv | 0.3339 | 1.3639 | 0.5040 | 0.4913 |
+| SQuAD | norm-cv | 0.3377 | 1.3866 | 0.5065 | 0.4940 |
+D.9 Negation Diagnostics and Qualitative Attention Illustration
+We include a controlled negation perturbation as a diagnostic of uncertainty sensitivity under semantic destabilization rather than as a correctness test.
+Because negation can legitimately change the true label, label-dependent quantities such as AccNeg and Err, both computed against the original labels, are only partially informative.
+We therefore emphasize label-free stability signals that reflect whether the model expresses increased uncertainty under the perturbation: the change in confidence (Conf), the change in the margin between the top-1 and top-2 predicted probabilities (Margin), and the change in predictive entropy (Entropy), where larger entropy and smaller margin indicate greater uncertainty.
+This diagnostic tests whether uncertainty estimates respond directionally to a controlled semantic change, rather than whether the model is invariant to arbitrary linguistic rephrasings.
+Accordingly, we do not claim that calibration behavior under negation generalizes to all paraphrases or stylistic variants.
+Instead, the template below provides a standardized perturbation that enables consistent comparison across methods under the same intervention.
+| Method | AccOrig | AccNeg | Conf | Margin | Entropy | Err |
+| Base (deterministic) | 0.705 | 0.495 | ||||
+| UAT-LITE (MC; Q-only default) | 0.690 | 0.495 | ||||
+| UAT-LITE + TS (MC-mean logits) | 0.690 | 0.495 |
+Across methods, the negation perturbation lowers confidence and margin while increasing entropy, indicating that it induces greater predictive uncertainty.
+Because the perturbation may change the underlying label, these results are best interpreted as an uncertainty-responsiveness check, that is, a test of directional sensitivity to semantic destabilization, rather than as a measure of robust accuracy under label-preserving transformations.
+To aid interpretability, Fig.
+7 visualizes a representative example of uncertainty-weighted attention under negation.
+The goal is to illustrate how uncertainty-guided gating can induce small, structured redistributions in internal information flow.
+We do not treat attention as a faithful explanation mechanism, and this evidence is not used as quantitative evidence.
+Because attention weights are normalized by softmax, absolute maps can appear visually similar; we therefore additionally show a difference heatmap to highlight directional shifts.
+D.10 Behavioral analysis by linguistic bucket (MedQA)
+To examine when uncertainty-weighted attention affects reliability, we stratify MedQA items into coarse linguistic buckets using simple cues, including negation markers, numeric content, biomedical terms, and hedging markers.
+Table 12 reports calibration and predictive performance within each bucket.
+Two patterns are consistent across buckets.
+First, UAT-LITE alone tends to improve accuracy relative to Base+TS (e.g., 0.24760.2614 on negation, 0.24180.2514 on numeric, 0.23690.2578 on biomedical, and 0.23440.2583 on hedges), but it does not improve in-domain calibration relative to TS, which remains the stronger post-hoc calibrator.
+Second, UAT-LITE+TS yields the lowest ECE in every bucket (0.0110 vs.
+0.0201 on negation; 0.0010 vs.
+0.0257 on numeric; 0.0075 vs.
+0.0302 on biomedical; 0.0080 vs.
+0.0331 on hedges), while preserving the accuracy gains introduced by UAT-LITE.
+This pattern supports a complementary interpretation: uncertainty-weighted attention improves internal evidence aggregation, whereas TS further corrects output-level confidence.
+Although coverage at the 0.9 threshold is zero for all methods in these buckets, ECE, AURC, and accuracy still reveal meaningful behavioral differences.
+Overall, this analysis suggests that the benefit of UAT-LITE is not confined to a single linguistic phenomenon, but appears across several medically relevant sources of ambiguity and variation.
+| Bucket | N | Method | ECE | Cov@0.9 | AURC | Acc |
+| all | 1273 | Base+TS | 0.0248 | 0.0000 | 0.7261 | 0.2427 |
+| UAT-LITE | 0.0809 | 0.0000 | 0.7478 | 0.2498 | ||
+| UAT-LITE+TS | 0.0006 | 0.0000 | 0.7493 | 0.2498 | ||
+| negation | 723 | Base+TS | 0.0201 | 0.0000 | 0.7213 | 0.2476 |
+| UAT-LITE | 0.0677 | 0.0000 | 0.7358 | 0.2614 | ||
+| UAT-LITE+TS | 0.0110 | 0.0000 | 0.7362 | 0.2614 | ||
+| numeric | 1245 | Base+TS | 0.0257 | 0.0000 | 0.7231 | 0.2418 |
+| UAT-LITE | 0.0793 | 0.0000 | 0.7441 | 0.2514 | ||
+| UAT-LITE+TS | 0.0010 | 0.0000 | 0.7455 | 0.2514 | ||
+| biomedical | 574 | Base+TS | 0.0302 | 0.0000 | 0.7259 | 0.2369 |
+| UAT-LITE | 0.0716 | 0.0000 | 0.7371 | 0.2578 | ||
+| UAT-LITE+TS | 0.0075 | 0.0000 | 0.7357 | 0.2578 | ||
+| hedges | 751 | Base+TS | 0.0331 | 0.0000 | 0.7471 | 0.2344 |
+| UAT-LITE | 0.0702 | 0.0000 | 0.7351 | 0.2583 | ||
+| UAT-LITE+TS | 0.0080 | 0.0000 | 0.7362 | 0.2583 |
+D.11 Sensitivity Analysis of and
+We examine the sensitivity of UAT-Lite to two inference-time hyperparameters: the uncertainty weighting coefficient and the Monte Carlo budget .
+We evaluate a representative grid with and on SQuAD 2.0, which is particularly sensitive to epistemic uncertainty due to the presence of both answerable and unanswerable questions.
 Similar stability trends were observed on MNLI and SST-2 and are omitted for brevity.
-We report ECE and predictive accuracy to assess both calibration quality and task performance.
-Accuracy is reported here solely to verify that calibration gains do not arise from performance degradation.
-Across all configurations, calibration performance remains stable, with ECE values confined to a narrow interval between 0.064 and 0.070 (mean , standard deviation ).
-No sharp sensitivity to either or is observed.
-Increasing the number of Monte Carlo samples yields only marginal calibration improvements beyond , indicating diminishing returns from additional stochastic passes.
-Predictive accuracy remains effectively unchanged across the grid, varying by less than 0.5% absolute, suggesting that uncertainty-weighted attention primarily affects calibration rather than accuracy.
-Overall, these results indicate that UAT-Lite does not rely on precise hyperparameter tuning to achieve effective calibration and is robust across a broad range of reasonable inference-time settings.
-This stability supports the practicality of the method in deployment scenarios where extensive hyperparameter search is undesirable.
-| ECE | Accuracy | ||
-| 0.1 | 3 | 0.0699 | 0.7075 |
-| 0.1 | 5 | 0.0676 | 0.7075 |
-| 0.1 | 10 | 0.0649 | 0.7109 |
-| 0.5 | 3 | 0.0703 | 0.7059 |
-| 0.5 | 5 | 0.0657 | 0.7096 |
-| 0.5 | 10 | 0.0682 | 0.7068 |
-| 1.0 | 3 | 0.0669 | 0.7090 |
-| 1.0 | 5 | 0.0655 | 0.7102 |
-| 1.0 | 10 | 0.0640 | 0.7110 |
-Summary. ECE mean , std ; min , max (range ).
-D.2 Clinical Domain Calibration
-Table 7 reports calibration and accuracy on two clinical benchmarks with distinct task structures and uncertainty profiles: MedQA and PubMedQA.
-We emphasize that these clinical datasets are used only to assess domain transfer of uncertainty-aware attention to medically specialized language distributions, and should not be interpreted as evidence of clinical decision reliability.
-In particular, what constitutes “good uncertainty” in clinical settings is often decision-dependent and may involve asymmetric error costs (e.g., prioritizing recall for critical conditions), which are not explicitly modeled in our evaluation.
-On MedQA, Deep Ensembles achieve the lowest ECE (0.007) and the highest accuracy (0.268), reflecting the benefit of model diversity in discrete, fact-centric multiple-choice medical question answering.
-In contrast, UAT-Lite exhibits a slightly higher ECE (0.040) with marginal accuracy improvement over the baseline.
-This behavior is expected, as MedQA primarily evaluates factual recall with limited contextual ambiguity, constraining the benefits of uncertainty-aware attention.
-Importantly, UAT-Lite achieves competitive performance without the substantial computational and storage overhead associated with training and maintaining multiple model replicas.
-A different trend emerges on PubMedQA.
-While Deep Ensembles substantially reduce ECE (0.051), this improvement coincides with a marked drop in accuracy (0.520), indicating overly conservative confidence estimates.
-In contrast, UAT-Lite attains the highest accuracy (0.640) while maintaining calibration comparable to the baseline.
-This suggests that integrating uncertainty directly into the attention mechanism can improve evidence aggregation in long, semantically dense biomedical abstracts, where uncertainty arises from contextual interpretation rather than discrete fact retrieval.
-Overall, these results highlight a trade-off between ensemble-based calibration and inference-time uncertainty integration.
-While ensembles excel in fact-centric clinical QA, UAT-Lite offers a more favorable accuracy–calibration balance on tasks characterized by semantic ambiguity and long-context reasoning, without incurring additional training or parameter costs.
-Accordingly, we treat the clinical-domain results as a domain-transfer generalization stress test rather than a primary claim about clinical deployment.
-| Dataset | Method | ECE | Accuracy |
-| MedQA | Baseline | 0.033 | 0.246 |
-| Temperature Scaling | 0.023 | 0.246 | |
-| Deep Ensemble (5) | 0.007 | 0.268 | |
-| UAT-Lite | 0.040 | 0.249 | |
-| PubMedQA | Baseline | 0.114 | 0.620 |
-| Temperature Scaling | 0.128 | 0.620 | |
-| Deep Ensemble (5) | 0.051 | 0.520 | |
-| UAT-Lite | 0.103 | 0.640 |
-D.3 Character-Level Adversarial Robustness
-Table 8 evaluates robustness under character-level perturbations on SST-2. We apply random 5% character swaps to simulate typographical errors and adversarial perturbations.
-Both baseline BERT and UAT-Lite exhibit substantial accuracy degradation under attack, with baseline accuracy dropping from 0.920 to 0.816 and UAT-Lite from 0.922 to 0.812.
-Calibration also degrades substantially, with ECE increasing from 0.062 to 0.133 for baseline and from 0.055 to 0.132 for UAT-Lite.
-While UAT-Lite maintains slightly better clean accuracy (0.922 vs.
+We report ECE and accuracy to assess calibration quality and task performance; accuracy is included to verify that calibration changes are not driven by degraded performance.
+Table 13 shows that calibration remains stable across the grid: ECE lies in a narrow interval (mean , std ).
+No sharp sensitivity to is observed; across blocks, ECE varies by at most .
+Within each , increasing tends to reduce ECE relative to the reference (ECE is consistently non-positive), but gains diminish beyond : moving from yields most of the improvement at a moderate cost increase (MC Cost ), while provides smaller additional benefits despite a cost.
+Accuracy is similarly robust: changes within each block are small (Acc absolute), indicating that uncertainty-weighted attention primarily modulates calibration rather than task accuracy.
+Overall, the method is insensitive to precise tuning of and achieves stable calibration over a broad range of inference-time settings, supporting practicality when extensive hyperparameter search is undesirable.
+ECE Acc ECE Acc MC Cost (vs.
+) (vs.
+) (vs.
+) 0.1 3 0.0699 0.7075 0.0000 0.0000 1.00 0.1 5 0.0676 0.7075 1.67 0.1 10 0.0649 0.7109 3.33 0.5 3 0.0703 0.7059 0.0000 0.0000 1.00 0.5 5 0.0657 0.7096 1.67 0.5 10 0.0682 0.7068 3.33 1.0 3 0.0669 0.7090 0.0000 0.0000 1.00 1.0 5 0.0655 0.7102 1.67 1.0 10 0.0640 0.7110 3.33
+Summary. ECE mean , std ; min , max (range ). ECE/Acc are computed within each block relative to ; MC Cost scales linearly with .
+D.12 Character-Level Adversarial Robustness
+Table 14 evaluates robustness under character-level perturbations on SST-2.
+We apply random 5% character swaps to simulate typographical noise and simple adversarial corruption.
+Both the baseline BERT model and UAT-Lite exhibit substantial accuracy degradation under attack, with accuracy dropping from 0.920 to 0.816 for the baseline and from 0.922 to 0.812 for UAT-Lite.
+Calibration also deteriorates markedly, with ECE increasing from 0.062 to 0.133 for the baseline and from 0.055 to 0.132 for UAT-Lite.
+Although UAT-Lite achieves slightly better clean accuracy (0.922 vs.
 0.920) and clean calibration (ECE 0.055 vs.
-0.062), the robustness under attack is comparable between methods.
-This indicates that uncertainty-aware attention improves calibration on clean data but provides limited additional protection against adversarial perturbations.
-These results are consistent with prior work showing that calibration methods address epistemic uncertainty but do not directly defend against adversarial attacks Goodfellow et al.
+0.062), its robustness under attack remains broadly comparable to that of the baseline.
+These results suggest that uncertainty-aware attention improves calibration under clean conditions but provides limited additional robustness to character-level perturbations.
+This pattern is consistent with prior work showing that calibration-oriented methods primarily address uncertainty estimation and do not, by themselves, constitute adversarial defenses Goodfellow et al.
 (2014).
-Future work should explore combining UAT-Lite with specialized adversarial defenses such as adversarial training or certified robustness techniques.
+An important direction for future work is to combine UAT-Lite with dedicated adversarial-defense strategies, such as adversarial training or certified robustness methods.
 | Attack | Method | Clean Acc | Attack Acc | Clean ECE | Attack ECE |
@@
 | Character Swap | UAT-Lite | 0.922 | 0.812 | 0.055 | 0.132 |
-D.4 Full Linguistic Uncertainty Analysis
-Table 9 provides a fine-grained linguistic analysis of how UAT-Lite responds to diverse sources of uncertainty, spanning lexical, syntactic, epistemic, and logical phenomena.
-For negation, UAT-Lite remains uniformly confident across affirmative, negated, and clinical negation cases, yielding zero estimated uncertainty.
-This behavior indicates negation invariance, which is desirable for explicit clinical negation (e.g., “not stable”) but also suggests limited sensitivity to polarity reversal in non-clinical contexts.
-This invariance is consistent with the attention analysis in Fig.˜3, which shows that uncertainty-aware attention redistributes weight conservatively toward negation and predicate structure without inducing elevated output-level uncertainty for clear clinical negation.
-In cases of lexical ambiguity, the framework exhibits mixed behavior.
-When sufficient contextual disambiguation is present (“bank account”), UAT-Lite appropriately increases uncertainty.
-However, for genuinely ambiguous expressions (“bank closed”), the model remains highly confident, indicating overconfidence under unresolved polysemy.
-Partial sensitivity is observed for polysemous terms such as “bat,” where uncertainty increases but does not fully reflect ambiguity.
-A similar pattern arises for epistemic hedges.
-Strong assertions (“definitely”) and weak hedges (“probably,” “possibly”) are treated with near-maximal confidence, suggesting hedge insensitivity.
-In contrast, explicit modal uncertainty (“might”) leads to a measurable increase in uncertainty, indicating that the model responds primarily to strong lexical uncertainty markers rather than gradated epistemic cues.
-For syntactic complexity, UAT-Lite demonstrates heightened sensitivity.
-Medium-depth syntactic structures produce the largest uncertainty increase, reflecting increased structural processing difficulty.
-Interestingly, deeper structures elicit a non-monotonic response, suggesting that the model may partially adapt to highly regular but complex constructions rather than uniformly increasing uncertainty with depth.
+D.13 Full Linguistic Uncertainty Analysis
+Table 15 provides a fine-grained analysis of how UAT-Lite responds to diverse sources of uncertainty spanning lexical, syntactic, epistemic, and logical phenomena.
+For negation, UAT-Lite remains uniformly confident across affirmative, negated, and clinical-negation cases, yielding near-zero estimated uncertainty.
+This behavior indicates invariance to explicit negation cues, which is desirable for clear clinical negation (e.g., “not stable”) but also suggests limited sensitivity to polarity reversal in more general contexts.
+This pattern is consistent with the attention analysis in Fig.˜7, which shows that uncertainty-aware attention redistributes weight conservatively toward negation and predicate structure without producing elevated output-level uncertainty for clear clinical negation.
+For lexical ambiguity, the behavior is mixed.
+When contextual disambiguation is available (“bank account”), UAT-Lite shows a meaningful increase in uncertainty.
+However, for genuinely ambiguous expressions (“bank closed”), the model remains highly confident, indicating residual overconfidence under unresolved polysemy.
+A similar intermediate pattern appears for polysemous triggers such as “bat,” where uncertainty increases but does not fully reflect the underlying ambiguity.
+A related pattern appears for epistemic hedges.
+Strong assertions (“definitely”) and weaker hedges (“probably,” “possibly”) are treated with near-maximal confidence, indicating limited sensitivity to graded epistemic marking.
+By contrast, explicit modal uncertainty (“might”) leads to a measurable increase in uncertainty, suggesting that the model responds more strongly to direct lexical uncertainty markers than to subtler hedging cues.
+For syntactic complexity, UAT-Lite shows clearer sensitivity.
+Medium-depth structures produce the largest increase in uncertainty, consistent with greater structural processing difficulty.
+Deeper structures, however, elicit a non-monotonic response, suggesting that the model may partially adapt to highly regular but more complex constructions rather than increasing uncertainty monotonically with depth.
 Finally, under logical contradictions, UAT-Lite appropriately increases uncertainty for inconsistent inputs (“healthy dying”) while maintaining high confidence for coherent statements.
-This indicates effective detection of semantic inconsistency, consistent with the framework’s uncertainty propagation across layers.
+This pattern indicates sensitivity to semantic inconsistency and is consistent with the framework’s uncertainty propagation across layers.
 | Phenomenon | Trigger | Base Conf | MB Conf | Unc | Diagnostic |
@@
 All models evaluated in this section are encoder-only, BERT-derived transformer architectures.
-Generalization to fundamentally different architectures (e.g., decoder-only models such as GPT or LLaMA, or state-space models) is outside the scope of the current study and is left for future work.
+Generalization to fundamentally different architectures, such as decoder-only models (e.g., GPT or LLaMA) or state-space models, is outside the scope of the current study and is left for future work.
 E.1 Generalization Across BERT Backbones
 Table 16 evaluates whether the calibration benefits of UAT-Lite generalize across BERT-based transformer backbones pretrained on heterogeneous corpora and applied to domain-specific benchmarks.
-The evaluated models span general-purpose (BERT-base, RoBERTa-base), clinical (ClinicalBERT), biomedical (BioBERT), and scientific (SciBERT) representations, thereby testing robustness across substantially different linguistic and semantic distributions.
-Across evaluated domains, UAT-Lite reduces ECE for most pretrained backbones, yielding an average relative improvement of approximately 32% (Table 10).
-The strongest gain is observed for BERT-base on SQuAD, where ECE decreases by 52% (0.130 0.062).
-This result is particularly significant because SQuAD contains both answerable and unanswerable questions, making it highly sensitive to epistemic uncertainty.
-The large improvement indicates that uncertainty-weighted attention enables the model to better recognize ambiguity during evidence aggregation rather than producing overconfident predictions.
-In domain-specialized settings, BioBERT on PubMedQA achieves a 40% reduction in ECE, suggesting that uncertainty-aware attention is especially beneficial for long, evidence-dense biomedical text that requires integrating information across multiple sentences.
-ClinicalBERT on MedQA shows a 25% improvement, indicating that the method remains effective even for fact-oriented medical question answering, although gains are smaller due to relatively strong baseline calibration.
-By contrast, RoBERTa-base on MNLI exhibits only a marginal 2% improvement.
-MNLI is a large, high-resource dataset with strong pretrained representations and comparatively well-calibrated baselines, leaving limited room for additional calibration gains.
-SciBERT on SQuAD similarly shows a modest 10% reduction, suggesting that pretraining on highly specialized scientific corpora may already encode task-relevant uncertainty patterns.
+The evaluated models span general-purpose (BERT-base, RoBERTa-base), clinical (ClinicalBERT), biomedical (BioBERT), and scientific (SciBERT) representations, thereby testing robustness across substantially different linguistic and semantic domains.
+Across the evaluated settings, UAT-Lite reduces ECE for most pretrained backbones, yielding an average relative improvement of approximately 32% (Table 16).
+The largest gain is observed for BERT-base on SQuAD, where ECE decreases by 52% (0.130 0.062).
+Because SQuAD contains both answerable and unanswerable questions, it places substantial demands on ambiguity recognition and confidence estimation.
+The large improvement in this setting suggests that uncertainty-weighted attention helps the model better calibrate predictions when evidence aggregation is uncertain.
+Domain-specialized models also benefit, although to different degrees.
+BioBERT on PubMedQA shows a 40% reduction in ECE, suggesting that uncertainty-aware attention is useful for evidence-dense biomedical text.
+ClinicalBERT on MedQA shows a 25% improvement, indicating that the method remains effective in fact-oriented medical question answering, although the gain is smaller than in some other settings.
+By contrast, RoBERTa-base on MNLI shows only a marginal 2% improvement.
+MNLI is a high-resource benchmark with strong pretrained representations and a relatively well-calibrated baseline, leaving less room for further gains.
+SciBERT on SQuAD similarly shows a modest 10% reduction, suggesting that pretraining on specialized scientific corpora may already capture some task-relevant uncertainty structure.
+Taken together, these results indicate that the calibration benefits of UAT-Lite transfer across multiple BERT-derived backbones and application domains, with the largest gains arising in settings characterized by greater ambiguity or less favorable baseline calibration.
 | Base Model | Domain | Dataset | Baseline ECE | +UAT-Lite ECE | Improvement |
@@
 E.2 Effect of Model Capacity on Calibration
-Table 11 analyzes the effect of model capacity on calibration improvement by applying UAT-Lite across different sizes of the BERT architecture, ranging from BERT-tiny (4.4M parameters) to BERT-large (335.1M parameters).
+Table 17 analyzes the effect of model capacity on calibration improvement by applying UAT-Lite across BERT models ranging from BERT-tiny (4.4M parameters) to BERT-large (335.1M parameters).
 Unlike Table 16, which varies domains and pretraining corpora, this analysis holds the architecture family fixed and isolates the role of parameter scale.
-Small and mid-sized models benefit substantially from uncertainty-weighted attention.
-BERT-mini achieves the largest relative improvement, with a 58% reduction in ECE, followed by BERT-base with an approximately 53% improvement (52–53% depending on rounding and binning resolution).
-These models exhibit sufficient representational capacity to express epistemic uncertainty, while still relying on attention mechanisms that benefit from explicit uncertainty modulation.
+Small and mid-sized models benefit most from uncertainty-weighted attention.
+BERT-mini achieves the largest relative improvement, with a 58% reduction in ECE, followed by BERT-base with an approximately 53% reduction.
+These models appear to have sufficient representational capacity to express epistemic uncertainty while still benefiting from explicit uncertainty modulation within attention.
 BERT-tiny and BERT-small show more modest improvements (12% and 15%, respectively).
-In very small models, limited representational capacity constrains the extent to which uncertainty-aware reasoning can reshape internal attention patterns, resulting in smaller but still consistent calibration gains.
-In contrast, BERT-large exhibits degraded calibration, with ECE increasing by 13% when UAT-Lite is applied.
-We hypothesize that in highly overparameterized models, stochastic perturbations introduced via Monte Carlo sampling may disrupt already well-formed attention patterns, leading to excessive variance propagation rather than meaningful uncertainty modeling.
-This observation suggests that uncertainty-weighted attention is most effective in the regime of small to mid-sized transformers, where model capacity and stochasticity appear to be better balanced.
-Latency measurements further contextualize this behavior.
-While inference-time overhead remains moderate for mid-sized models (e.g., 7 ms 39 ms for BERT-base), it increases substantially for larger architectures (26 ms 134 ms for BERT-large), reinforcing the trade-off between calibration gains and computational cost.
+For very small models, limited representational capacity may constrain the extent to which uncertainty-aware inference can reshape internal attention patterns, yielding smaller but still consistent gains.
+In contrast, BERT-large shows degraded calibration, with ECE increasing by 13% after applying UAT-Lite.
+One possible explanation is that, in highly overparameterized models, stochastic perturbations introduced through Monte Carlo sampling may disrupt already stable attention patterns, leading to excess variance propagation rather than improved uncertainty estimation.
+Latency measurements further contextualize these results.
+Although inference-time overhead remains moderate for mid-sized models (e.g., 7 ms 39 ms for BERT-base), it increases substantially for larger architectures (26 ms 134 ms for BERT-large).
+Overall, this analysis suggests that uncertainty-weighted attention is most effective in the small-to-mid-sized regime, where model capacity and stochasticity appear to be better balanced.
 | Model | Params | Baseline ECE | UAT-Lite ECE | Rel. | Latency |
@@
 We first analyze its asymptotic and practical computational complexity relative to standard transformer inference and ensemble-based uncertainty methods.
-We then quantify inference-time overhead introduced by Monte Carlo sampling and describe the experimental setup, hardware configuration, and reproducibility details required to replicate all reported results.
-Throughout, we emphasize that UAT-Lite operates strictly at inference time and does not modify pretrained parameters or training cost.
+We then quantify the inference-time overhead introduced by Monte Carlo sampling and summarize the experimental setup, hardware configuration, and reproducibility details required to replicate the reported results.
+Throughout, we emphasize that UAT-Lite operates strictly at inference time and does not alter pretrained parameters or training cost.
 F.1 Computational Complexity Analysis
 UAT-Lite preserves the parameter count and training cost of the base transformer.
-All uncertainty-aware mechanisms operate strictly at inference time and do not introduce additional trainable parameters or modify learned weights.
-Additional computational cost arises solely from Monte Carlo (MC) sampling during inference:
+All uncertainty-aware mechanisms operate strictly at inference time and introduce no additional trainable parameters or architectural modifications.
+The additional cost arises solely from Monte Carlo (MC) sampling during inference:
 Parameters: identical to the base model (e.g., 109M for BERT-base)
@@
 Calibration cost: one scalar temperature per task (optional)
-Inference cost: stochastic forward passes (typically )
-From a complexity perspective, inference-time cost scales linearly with the number of Monte Carlo samples .
-In contrast to ensemble-based approaches, which require training, storing, and evaluating multiple independent models, UAT-Lite incurs no additional training or memory overhead.
-As a result, its total computational footprint remains substantially lower than Deep Ensembles for comparable calibration gains.
+Inference cost: stochastic forward passes ()
+Inference-time complexity therefore scales linearly with the MC budget .
+Table 18 reports a canonical end-to-end GPU benchmark at under a sequential MC evaluation loop.
+Relative to deterministic inference, UAT-Lite incurs approximately a slowdown, whereas temperature scaling adds negligible overhead because it applies only a scalar rescaling to the logits.
+Unlike Deep Ensembles, which require training, storing, and evaluating multiple independent models, UAT-Lite introduces no additional training or model-storage cost, making it a lighter-weight alternative when retraining is undesirable.
+| Method | Latency (ms) | Ratio | |
+| Base (det.) | 62.93 | 0.29 | 1.00 |
+| Base + TS | 63.13 | 0.27 | 1.00 |
+| UAT-Lite (M=10) | 1426.90 | 78.87 | 22.68 |
+| UAT-Lite + TS (M=10) | 1486.61 | 7.26 | 23.62 |
 F.2 Inference-Time Overhead
@@
 When executed sequentially, runtime increases approximately linearly with .
-For BERT-base, this corresponds to a measured slowdown of approximately relative to deterministic inference when using .
-Despite this overhead, inference-time cost remains practical for many deployment settings, particularly in scenarios where reliability and calibrated confidence are prioritized over raw throughput.
-Importantly, this cost is incurred only when uncertainty-aware inference is enabled and can be adjusted dynamically by reducing for high-confidence inputs, as discussed in Section Limitations.
+Observed slowdowns depend on the Monte Carlo budget and the latency measurement protocol, including batching, sequence length, and whether MC passes are executed sequentially or vectorized.
+Throughout this work, we report latency under the canonical benchmark described in F.4, under which overhead scales approximately linearly with for sequential MC inference.
+Despite this overhead, inference-time cost remains practical in settings where reliability and calibrated confidence are prioritized over raw throughput.
+Importantly, this cost is incurred only when uncertainty-aware inference is enabled and can in principle be reduced by lowering in settings where a smaller stochastic budget is acceptable, as discussed in Section Limitations.
 F.3 Reproducible Experimental Setup
@@
 Learning rate:
-Learning rate schedule: linear warmup for first 500 steps, followed by linear decay
+Learning rate schedule: linear warmup for the first 500 steps, followed by linear decay
 Batch size: 16
@@
 UAT-Lite configuration.
-Inference via Monte Carlo sampling with stochastic forward passes
+Inference via Monte Carlo sampling with stochastic forward passes (default for the main revision tables; is additionally reported where noted)
 Component-specific dropout rates:
@@
 Uncertainty weighting coefficient: (set once on development data and held fixed across tasks)
-Sensitivity analysis for and reported in §˜D.1
+Sensitivity analysis for and reported in D.11
 F.4 Hardware, Latency, and Code Availability
@@
 All experiments are conducted on a single NVIDIA A100 GPU (40GB).
-Unless otherwise stated, latency is measured as end-to-end forward-pass wall-clock time per example, averaged over 200 runs after 50 warm-up iterations.
-We use fixed sequence lengths (128 for SST-2 and MNLI; task-standard truncation for SQuAD answerability) and batch size 1.
-For stochastic methods, latency is reported under the same Monte Carlo budget used during evaluation.
-Reported slowdowns reflect the ratio between stochastic inference and deterministic baseline inference under identical conditions.
+Unless otherwise stated, latency is measured as end-to-end forward-pass wall-clock time, averaged over 200 runs after 50 warm-up iterations.
+We report a single canonical latency benchmark throughout the paper: batch size , sequence length , a sequential MC loop without vectorization across MC samples, and a single GPU.
+Latency is reported per batch and expressed as a ratio relative to deterministic inference under the same conditions.
+For stochastic methods, latency is measured using the same Monte Carlo budget used during evaluation.
+Reported slowdowns therefore reflect the ratio between stochastic and deterministic inference under identical measurement settings.
 Code availability.
-To ensure reproducibility during the anonymous review process, we provide an anonymized repository containing the complete implementation, including training scripts, evaluation pipelines, and configuration files.
+To support reproducibility during anonymous review, we provide an anonymized repository containing the complete implementation, including training scripts, evaluation pipelines, and configuration files.
 All experiments are implemented using HuggingFace Transformers, PyTorch, and Python.